### v1

# Token Preference Optimization with Self-Calibrated Visual-Anchored Rewards for Hallucination Mitigation

Abstract
Direct Preference Optimization (DPO) has been demonstrated to be highly effective in mitigating hallucinations in Large Vision Language Models (LVLMs) by aligning their outputs more closely with human preferences. Despite the recent progress, existing methods suffer from two drawbacks: 1) Lack of scalable token-level rewards; and 2) Neglect of visual-anchored tokens. To this end, we propose a novel Token Preference Optimization model with self-calibrated rewards (dubbed as TPO), which adaptively attends to visual-correlated tokens without fine-grained annotations. Specifically, we introduce a token-level visual-anchored reward as the difference of the logistic distributions of generated tokens conditioned on the raw image and the corrupted one. In addition, to highlight the informative visual-anchored tokens, a visual-aware training objective is proposed to enhance more accurate token-level optimization. Extensive experimental results have manifested the state-of-the-art performance of the proposed TPO. For example, by building on top of LLAVA-1.5-7B, our TPO boosts the performance absolute improvement for hallucination benchmarks.
Token Preference Optimization with Self-Calibrated Visual-Anchored Rewards for Hallucination Mitigation
Jihao Guthanks: Equal contribution.thanks: Work done during an internship at Alibaba Group.1, Yingyao Wang11footnotemark: 11, Meng Cao2, Pi Bu1, Jun Songthanks: Corresponding Author.1, Yancheng He1, Shilong Li1, Bo Zheng1 1Alibaba Group 2Mohamed bin Zayed University of Artificial Intelligence {gujihao.gjh, wangyingyao.wyy, jsong.sj}@taobao.com
1 Introduction
Recently, Large Vision Language Models (LVLMs) have showcased their remarkable capabilities in handling multimodal information, excelling in tasks such as image captioning, visual question-answering, and complex visual reasoning Team et al. (2023); Bai et al. (2023); Hurst et al. (2024); Yang et al. (2023). Specifically, by integrating pre-trained language models with meticulously designed visual encoders, LVLMs are capable of effectively capturing the semantic correlations between visual and textual data. This integration supports more accurate and contextually relevant tasks of visual understanding and generation.
Figure 1: An example of visual Q&A. The upper box contains the ground truth answer, while the lower box shows the LVLM responses before and after training with our method. In each box, we visualize the rewards for each token which can reflect the degree of visual anchoring, with the top representing scores before training and the bottom after. Scoring is detailed in Equation 5, and we’ve applied sigmoid normalization in this score.
Despite the advancements, the issue of “hallucination”, where the generated responses are not grounded in the input visual contexts, greatly impedes the reliability and practical deployment of LVLMs Liu et al. (2024a); Bai et al. (2024). To alleviate this, various methods have been proposed from the perspectives of data quality Liu et al. (2023); Zhai et al. (2023) and inference-time strategies Yin et al. (2023); Zhou et al. (2023); Huang et al. (2024). Recently, direct preference optimization (DPO) Rafailov et al. (2024) is introduced to align outputs with human preferences, therefore reducing the risk of generating hallucinatory or nonsensical responses.
Existing DPO-like methods, however, still suffer from two drawbacks: 1) Lack of scalable token-level rewards. The fine-grained token-level rewards enable precise adjustments to individual parts of generated responses. Existing methods, however, either provide global sentence rewards or rely on manual efforts for fine-grained segment-level annotations Yu et al. (2024b). Therefore, designing a scalable token-level reward generation strategy has become a clearly defined necessity (c.f. Table 1); 2) Neglect of visual-anchored tokens: By “visual-anchored tokens”, we refer to response tokens that are essential and highly correlated with the input visual embeddings. RLHF-V assigns all the hallucinated segments with a fixed reward value. Recent studies Guan et al. (2024) attribute the hallucination issue to an inherent imbalance between the visual and textual modalities. Specifically, due to the large-scale pre-trained textual corpus, LVLMs tend to prioritize language-based information even at the costs of overriding the provided visual content. Therefore, we argue that not all the tokens are equal, i.e., visual-anchored tokens (e.g., glass in Figure 1) are more prone to hallucination and deserve great emphasis. As shown in Table 1, the concurrent pre-print V-DPO Xie et al. (2024) also focuses on visual-anchored tokens; however, it requires the additional construction of a synthetic dataset, whereas our method eliminates the need for any extra annotations.
To alleviate these aforementioned problems, we propose a novel Token Preference Optimization with self-calibrated rewards (dubbed as TPO), which rectifies the fine-grained token-level hallucinations and attends to visual-anchored tokens without the need of fine-grained annotations. Specifically, to mine the visual-anchored tokens, we compute the differences between the logits distributions of generated tokens conditioned on the raw image and the corrupted one. We regard this distribution difference as token-wise rewards. In Figure 1, we apply this visual-anchored score mining strategy on both golden truth and the generated responses. As shown, this strategy effectively helps highlight visual-anchored tokens. Then, we propose a token preference optimization loss by integrating the self-calibrated rewards into the vanilla DPO. In particular, we multiply the like-hood distribution with token-wise rewards to generate our desired visual-correlated ones.
Overall, the main contributions of this work are:
  • We propose TPO for hallucination mitigation in LVLMs, which implements token-level distribution rectification without the reliance of fine-grained manual annotations.
  • We mine visual-anchored tokens by comparing the response distributions conditioned on the raw image and the corrupted one.
  • Extensive experiments on the popular hallucination benchmarks demonstrate the state-of-the-art performance of the proposed TPO.
Methods Visual- Anchored Token- level Fine-grained Annotations
DPO
POVID
CSR
RLHF-V
V-DPO
TPO (Ours)
Table 1: Comparisons with hallucination mitigation methods from the perspective of whether attending to vision-anchored tokens, whether generating token-level rewards and whether requiring fine-grained annotations. The compared methods include DPO Rafailov et al. (2024), POVID Zhou et al. (2024a), CSR Zhou et al. (2024b), RLHF-V Yu et al. (2024b), V-DPO Xie et al. (2024) and our proposed TPO.
2 Related Works
2.1 LVLMs’ Hallucination
Leveraging the rich knowledge in large language models and the vision understanding capabilities of vision encoders, LVLMs have shown exceptional performance in image understanding and generation tasks Li et al. (2023b); Zhu et al. (2023). However, imbalances in parameters and data scale during pre-training can lead to LVLMs being overly influenced by biases in the language model, resulting in inadequate attention to visual information and potential hallucination issues Zhou et al. (2023); Zhang et al. (2024). Consequently, addressing the issue of hallucinations in LVLMs has become one of the key research focuses in this field.
Previous studies have mitigated hallucinations by enhancing training data quality, refining decoding strategies, and post-processing generated responses Huang et al. (2024); Leng et al. (2024); Yu et al. (2024a); Han et al. (2024); Chen et al. (2024); Zhou et al. (2023); Yin et al. (2023); Lee et al. (2023); Xia et al. (2024); Shao et al. (2024). While these methods can lead to more accurate responses, they do not fundamentally resolve the issue of inadequate visual information association in LVLMs.
2.2 Preference Learning Methods
More recently, reinforcement learning from human feedback (RLHF) Sun et al. (2023) is gradually becoming a prevalent approach to mitigate the hallucination. As a more direct and effective method, DPO Rafailov et al. (2024) and its variants are more widely utilized for preference alignment.
Several studies based on DPO focus on developing more robustly constructed preference data. For example, the POVID Zhou et al. (2024a) method constructs negative samples for preferred data by adding noise to the image and providing hallucinated patterns to guide the model to generate hallucinated responses. The pre-structured data will be employed for off-policy DPO training. Apart from these works, RLAIF Yu et al. (2024c) and CSR Zhou et al. (2024b) methods, which are built upon on-policy DPO strategy, construct preference pairs by iteratively performing self-rewarding to select preference pairs. A notable commonality among the aforementioned studies is that assigning response-level rewards for each generated sequence is insufficient for effectively aligning with genuinely hallucination-prone contents.
Other studies, RLHF-V Yu et al. (2024b) and V-DPO Xie et al. (2024), investigated this issue and achieved more fine-grained alignment of preference data. Nevertheless, this approach depends on resource-intensive annotations or data constructions and applies a fixed reward to all hallucinated segments, thus failing to account for the differing levels of relevance these segments may have to visual information. It is worth mentioning that CSR also considered this problem and introduced CLIP Radford et al. (2021) to calculate the relevance score between generated text and vision information as an additional reward. However, this method requires the introduction of an additional model, which reduces the training efficiency.
In this paper, we propose a token-level preference optimization method with self-calibrated visual-anchored rewards (TPO), aimed at addressing the aforementioned challenges. TPO facilitates finer-grained alignment in LVLMs, enhancing accuracy in visual information correlation and reducing hallucinations during response generation.
3 Methodology
Figure 2: Outline of our TPO pipline. The process is divided into three parts for each data at every training step. First, 1) add noise to the image, then, 2) calculate Self-Calibrated Visual-Anchored Rewards, and finally 3) perform Token Preference Optimization. At the end of each training step, we calibrate the model and calculate new Visual-Anchored Rewards for the next step.
3.1 Motivation
As illustrated in Figure 1, the generation of each token requires differing levels of integration with visual information. Tokens that are closely correlated with visual inputs are designated as "visual-anchored tokens" in this paper. It is observed that the generation of such visual-anchored tokens can be more susceptible to hallucinations. Furthermore, directly assigning rewards to entire generated sequences can lead to tokens independent of visual information consuming undue optimization resources. This misallocation can reduce the supervisory focus on visual-anchored tokens. To enhance training efficiency and ensure the model attends to relevant visual features, it is crucial to assign token-level rewards based on the reliance on visual information during training.
To achieve these goals, the proposed TPO mitigates hallucination from the following three key aspects: 1) scalable token-level rewards to optimize alignment; 2) automatic identification of visual-anchored tokens by the model itself; 3) self-calibration of reward signals based on visual correlation.
3.2 Preliminaries
DPO Rafailov et al. (2024) is designed to directly maximize the reward margin between positive and negative responses to align human preferences. Specifically, given a textual input $x$, a visual input $v$, a negative response $y_{l}$, and a preferred positive response $y_{w}$, the reward function is defined as $r(x,v,y_{l}/y_{w})$. According to the Bradley-Terry Bradley and Terry (1952) model, the probability of each response is as follows:
$p(y_{w}>y_{l})=\sigma(r(x,v,y_{w})-r(x,v,y_{l}))$ (1)
where the sigmoid function is denoted as $\sigma(\cdot)$.
In this paper, we focus on the off-policy DPO method, where the probability of preference data is derived from the optimal policy rather than the reward model. The reward function is designed as:
$r(x,v,y)=\beta\log\frac{\pi_{\theta}(y|x,v)}{\pi_{\text{ref}}(y|x,v)}$ (2)
In Equation (2), assuming a response $y\in\{y_{w},y_{l}\}$ has $N$ tokens $(y_{1},...,y_{i},...,y_{N})$, the output distribution of the current policy model $\pi_{\theta}(y|x,v)$ and the reference model $\pi_{\text{ref}}{(y|x,v)}$ can be expressed as $\prod_{y_{i}}p(y_{i}|x,v)$. On this basis, the formulation of a maximum likelihood objective is defined as:
$\mathcal{L}_{DPO}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,v,y_{w},y_{l})\sim D}\Bigg{[}\log\sigma$ (3)
$\Bigg{(}\beta\log\frac{\pi_{\theta}(y_{w}|x,v)}{\pi_{\text{ref}}(y_{w}|x,v)}-\beta\log\frac{\pi_{\theta}(y_{l}|x,v)}{\pi_{\text{ref}}(y_{l}|x,v)}\Bigg{)}\Bigg{]}$
It is demonstrated that DPO assigns equal rewards to all tokens in the responses. When applied to LVLMs for mitigating hallucinations, it lacks emphasis on visual-anchored tokens. Therefore, we propose a novel approach named TPO to mainly solve this problem.
3.3 TPO
3.3.1 Visual-Anchored Rewards
We propose a self-calibrated visual-anchored reward generation mechanism based on Equation (2). On the one hand, it can automatically identify visual-anchored tokens in the response during training. On the other hand, it adaptively and iteratively assigns different rewards for tokens according to their visual reliance. The proposed rewards can not only strengthen the attention to visual information for token generation but also align preferences.
To construct the visual-anchored rewards, we first introduce a score $s$ to measure the visual reliance degree of each token. For this purpose, we add noise into the embedding of the input image $v$ to obtain the corrupted image $v_{c}$:
$v_{c}(k)=\sqrt{\bar{\xi}_{k}}\cdot v+\sqrt{1-\bar{\xi}_{k}}\cdot\epsilon$ (4)
Here, we mainly refer Zhou et al. (2024a) and set the noise steps $k=500$, $\bar{\xi}_{k}=\prod_{i=0}^{k}\xi_{i}$, $\xi$ is a predefined noise parameter derived from a list with 1,000 equally spaced elements. [^1] More details can be found in Appendix A. Subsequently, we capture the changes in the probability distribution of each generated token before and after adding noise. For token $y_{i}$ of the response $y$, the calculation of $s_{y_{i}}$ is formulated as Equation (5). Notably, here $p_{log}$ refers to the raw logits output of the model, before applying the softmax transformation.
$s_{y_{i}}=p_{log}(y_{i}|x,v,y_{<i})-p_{log}(y_{i}|x,v_{c},y_{<i})$ (5)
We have shown an example of the distribution of $s$ before and after training in Figure 1, it is shown that visual-anchored tokens will obtain higher scores. In other words, $s$ has the ability to reflect the visual relevance of each token $y_{i}$, this also lays the foundation for the ability of our visual-anchored rewards to self-calibrate based on the visual correlation of tokens.
Furthermore, we introduce a self-calibration process to generate the final visual-anchored rewards. This process aims to ensure that positive samples receive higher rewards than negative samples while optimizing the visual relevance of visual-anchored tokens in all responses. Formally, the process is defined as follows, where the sigmoid function $\text{sigmoid}(\cdot)$ is used for normalization.
$c_{y_{i}}=\begin{cases}a+\text{sigmoid}(s_{y_{i}})&\text{if }y_{i}\in y_{w}\\ a+1-\text{sigmoid}(s_{y_{i}})&\text{if }y_{i}\in y_{l}\end{cases}$ (6)
The calculated $c_{y_{i}}$ is adopted as the final reward in TPO, where $c_{y_{i}}\in(a,a+1)$. In this way, $c$ and $s$ are positively correlated for positive examples, whereas for negative ones, the opposite holds true. In particular, $p(y_{i}|x,v,y_{<i})=p(y_{i}|x,v_{c},y_{<i})$, means that the token does not associate any visual information. In this case, no visual-anchored reward should be assigned for this token. To this end, we set $a=0.5$ in Equation (6), so that when $s=0$, $c=1$, the rewards will not take effect.
3.3.2 Token Preference Optimization
After obtaining the reward $c_{y_{i}}$ to $y_{i}$, the output cumulative distribution can be calculated:
$\pi(y|x,v)=\prod_{y_{i}\in\mathcal{Y}}c_{y_{i}}$ (7)
Especially, when $c_{y_{i}}=1$, the probability of $y_{i}$ will not be accumulated. By multiplying the probability distribution with the visual-anchored rewards, we obtain a novel KL-constrained reward maximization objective:
$\max_{\pi}E_{(x,v,y)}\Bigg{[}\,r^{\prime}(x,v,y)-\beta D_{KL}\Bigg{(}\left[\pi_{\theta}(y|x,v)\right.$ (8)
×$\left.\times\pi_{\theta}^{v}(y|x,v)\right]\Bigg{\|}\left[\pi_{\text{ref}}(y|x,v)\times\pi_{\text{ref}}^{v}(y|x,v)\right]\Bigg{)}\Bigg{]}$
Thus, the optimal solution formula for the maximization objective of the KL-constrained reward is as follows:
$\pi_{\theta}(y|x,v)\times\pi_{\theta}^{v}(y|x,v)=\frac{1}{Z(x)}\left[\pi_{\text{ref}}(y|x,v)\times\right.$ (9)
$\left.\pi_{\text{ref}}^{v}(y|x,v)\right]\exp\left(\frac{1}{\beta}r^{\prime}(x,v,y)\right)$
Rearranging the Eq (9), we obtain the new reward function:
$r^{\prime}(x,v,y)$ (10)
$=\beta\log\frac{\pi_{\theta}(y|x,v)\times\pi_{\theta}^{v}(y|x,v)}{\pi_{\text{ref}}(y|x,v)\times\pi_{\text{ref}}^{v}(y|x,v)}+\beta Z(x)$
$=\beta\sum_{y_{i}\in\mathcal{Y}}\left[\log\left(p_{\theta}(y_{i}|x,v,y_{<i})\cdot c_{\theta}\right)\right.$
$\quad\left.-\log\left(p_{\text{ref}}(y_{i}|x,v,y_{<i})\cdot c_{\text{ref}}\right)\right]+\beta Z(x)$
$=\beta\sum_{y_{i}\in\mathcal{Y}}\left[\log p_{\theta}(y_{i}|x,v,y_{<i})-\right.$
$\quad\left.\log p_{\text{ref}}(y_{i}|x,v,y_{<i})+\log\frac{c_{\theta}}{c_{\text{ref}}}\right]+\beta Z(x)$
Compared to the original reward function in DPO (Equation (2)), we multiply each $p(y_{i}|x,v,y_{<i})$ by the generated visual-anchored rewards $c_{y_{i}}$ at the token level. For $c_{\theta}$, its computation is continuously updated at each step during training as the model changes. In the calculation of each token in the entire reward function, we add a term $log\frac{c_{\theta}}{c_{\text{ref}}}\in(-log3,log3)$, which has a reasonable upper and lower bound. For positive samples, this term is expected to increase, while for negative samples, it is expected to decrease. Due to the different methods of calculating $c_{y_{i}}$ that we set in Equation (6), this will encourage the increase of $s_{y_{i}}$ during the training process, making the token generation focus more on visual information.
Thus, following the Bradley-Terry model, when given the positive and negative samples $\mathcal{D}=\{x^{(i)},v^{(i)},y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{N}$, we obtain our maximum likelihood objective:
$\mathcal{L}_{TPO}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,v,y_{w},y_{l})\sim D}\bigg{[}\log\sigma$ (11)
$\quad\Bigg{(}\beta\log\frac{\pi_{\theta}(y_{w}\big{|}x,v)\times\pi_{\theta}^{v}(y_{w}\big{|}x,v)}{\pi_{\text{ref}}(y_{w}\big{|}x,v)\times\pi_{\text{ref}}^{v}(y_{w}\big{|}x,v)}-$
$\quad\beta\log\frac{\pi_{\theta}(y_{l}\big{|}x,v)\times\pi_{\theta}^{v}(y_{l}\big{|}x,v)}{\pi_{\text{ref}}(y_{l}\big{|}x,v)\times\pi_{\text{ref}}^{v}(y_{l}\big{|}x,v)}\Bigg{)}\bigg{]}$
Method AMBER MMHal HallusionBench General Benchmarks
Acc F1 Score Hall $\downarrow$ Easy Hard aAcc SEED MMB LLaVA MM-Vet
LLAVA-1.5-7B 71.7 74.3 2.01 61.46 42.64 41.16 47.21 66.1 73.3 65.6 31.6
+ DPO 77.5 82.1 2.14 58.33 37.36 37.21 43.84 66.4 73.3 69.1 31.6
+ CSR 73.2 76.1 2.05 60.42 43.08 41.16 47.48 65.9 73.0 68.9 31.0
+ POVID 71.9 74.7 2.26 55.21 42.86 41.63 47.56 66.1 73.2 68.2 31.7
+ RLHF-V 74.8 78.5 2.02 60.42 42.20 43.72 48.27 66.1 73.1 68.0 32.3
+ V-DPO 81.6 2.16 56.00 51.63
+ TPO (Ours) 79.3 85.0 2.47 51.04 41.76 48.37 50.22 66.6 73.6 70.2 33.0
LLAVA-1.5-13B 71.3 73.1 2.38 53.13 44.40 36.51 46.94 68.2 76.7 73.1 36.1
+ DPO 83.2 86.9 2.47 51.04 45.49 43.49 50.22 68.6 76.6 72.8 37.5
+ RLHF-V 79.2 82.3 2.50 52.08 43.96 40.00 48.27 68.2 76.7 76.7 38.5
+ TPO (Ours) 83.9 88.0 2.72 45.83 44.40 46.05 50.93 68.7 76.8 72.8 36.2
Table 2: Performence of LLaVA-1.5 on hallucination and general benchmarks. Score and Hall refer to the overall GPT-4 score and hallucination rate, respectively. Easy represents the accuracy of with original images, hard represents the accuracy with manually edited challenging images, and aAcc is the average accuracy for each question. The results for POVID Zhou et al. (2024a) and CSR Zhou et al. (2024b) are based on our testing of their open-source model weights, while the results for V-DPO Xie et al. (2024) are taken from previous work. (bold: the best score; underline: the second best)
4 Experiment
4.1 Setup
Aligning with previous DPO-based approaches on hallucination mitigation, we adopt the popular LVLM, LLaVA-1.5 Liu et al. (2024b), as the backbone model to validate the effectiveness of our TPO. As for the dataset, we directly utilize the preference pairs provided by RLHF-V (5k) without their fine-grained human annotations.
Benchmarks: We primarily conduct the experiments on three hallucination benchmarks: (1) AMBER Wang et al. (2023): a multi-dimensional hallucination benchmark with over 15k samples. We mainly focus on its discriminative part and report Accuracy and F1 metrics referencing Yu et al. (2024c); (2) MMHal-Bench Sun et al. (2023): it measures the hallucination rate and informativeness of responses. (3) HallusionBench Guan et al. (2024): it evaluates visual illusions and knowledge hallucinations through systematically structured discriminative tasks. We also test the performance on four general evaluation benchmarks: (1) SEED Bench Li et al. (2023a): a benchmark for LVLMs on generative comprehension. (2) MMBench Liu et al. (2025): a comprehensive benchmark designed to evaluate the capabilities across various tasks and modalities. (3) LLaVA Bench Liu et al. (2024c): a benchmark for evaluating multi-modal conversation, detailed description, and complex reasoning. (4) MM-Vet Yu et al. (2023): a benchmark assessing for integrated capabilities.
Baselines: In this work, we mainly compare TPO with the SFT model of LLaVA-1.5-7B, as well as with DPO and V-DPO Xie et al. (2024) methods trained using RLHF-V Yu et al. (2024b) data, along with two improved methods, CSR Zhou et al. (2024b) and POVID Zhou et al. (2024a). Moreover, to evaluate the effectiveness and robustness of TPO as the model size increases, we further compare it with the LLaVA-1.5-13B SFT model, and also against DPO and RLHF-V methods.
Figure 3: Performance curves with the change of the noise steps-(a) and the change of parameter $a$-(b), We separately present the F1 of AMBER, the hallucination rate of MMHAL, the aACC of HallusionBench, and the average value of the general benchmarks. More detailed metrics can be found in the Appendix B.
4.2 Main Results
[…]
Notably, on the HallusionBench evaluation metrics, "Easy" represents the accuracy of original image-based questions, which tend to rely on prior knowledge, while "Hard" represents the accuracy of questions based on manually edited images, which tend to rely on visual information. Our method leads to the most significant improvement for the original model on hard questions, with a slight improvement on easy questions. This indicates that, compared to other methods, our approach enables the model to focus more on visual information rather than the textual prior knowledge to provide accurate answers.
On general benchmarks, compared to other methods, our approach remains stable against the original LLaVA model and achieves the greatest improvement on most benchmarks. We attribute it to that our method helps the model associate with more visual information when answering questions. This shows that our approach can improve hallucination issues while maintaining good performance on general evaluation tasks.
Method AMBER MMHal HallusionBench General Benchmarks
Acc F1 Score Hall $\downarrow$ Easy Hard aAcc SEED MMB LLaVA MM-Vet
LLAVA-1.5-7B 71.70 74.3 2.01 61.46 42.64 41.16 47.21 66.1 73.3 65.6 31.6
Only Win 79.10 84.5 2.24 56.25 44.62 46.05 50.40 66.6 73.6 69.8 31.7
Only Loss 79.20 84.8 2.33 53.13 42.20 47.91 49.87 66.6 73.5 70.7 32.0
Opposite 75.30 80.7 1.91 64.58 42.42 45.58 48.63 65.6 73.1 68.9 32.1
TPO (Ours) 79.30 85.0 2.47 51.04 41.76 48.37 50.22 66.6 73.6 70.2 33.0
Table 3: Ablation Studies. Performence of LLaVA-1.5 on hallucination and general benchmarks. (bold: the best score)
Figure 4: Comparison of attention weights for LLaVA-1.5 before and after TPO training. Each horizontal line represents the mean of that data set. The blue section on the left response incorrectly, with many ’visual-anchored tokens’ tokens having high attention weights but resulting in hallucinated responses (e.g. USB, cloud). The red section on the right answered correctly.
4.3 Ablation Studies
To further demonstrate how our method helps improve model performance by reducing hallucinations, we conduct ablation experiments for the noisy steps, the introduced parameter $a$, and the self-calibrated process for rewards. These experiments are performed on LLaVA-1.5-7B consistently with the main experiments, including hallucination and general evaluation benchmarks.
Noise Steps : We ablate on the noise steps in Figure 3 (a). As shown, the optimal performance is achieved at the step of 500. This medium corruption enables the model to grasp the general outline of the image while missing the detailed contents, which is prone to generate hallucinations of the visual-anchored tokens.
Parameter $a$ : We conduct experiments by varying the parameter $a$ introduced in Equation (6) with the results shown in Figure 3 (b). By setting $a=[0,0.5,1]$, we observed consistently good performance across all configurations. This suggests that effective performance is achieved as long as the reward mechanism successfully highlights token differences and identifies visually anchored tokens. Notably, the best overall results are obtained with $a=0.5$, validating our proposed method and hypothesis. This indicates that when the visual-anchored score $s=0$, setting $c=1$, not introducing additional reward signals can yield better outcomes.
Visual-Anchored Rewards : Furthermore, Table 3 demonstrates that TPO can enhance model performance when rewards are assigned separately to positive and negative samples, achieving results comparable to those obtained by rewarding both simultaneously. However, by providing opposite rewards to positive and negative samples, where rewards are negatively correlated with the visual relevance of positive samples and positively correlated with that of negative samples, TPO’s performance significantly deteriorates. In some metrics, this approach yields even poorer results than the original LLaVA-1.5 model. This further underscores the validity of the designation of visual-anchored rewards.
By integrating these selection methods, we achieve superior overall performance, thereby highlighting the value and efficacy of our approach. This integration underscores both the simplicity and effectiveness of our method in achieving robust results.
4.4 Analysis
Attentions : To further substantiate the effectiveness of our method in enhancing visual alignment, we visualize the relevance between response and image tokens, both before and after training with TPO. In this statistic, the relevance is quantified using the sum of attention weights. As depicted in Figure 4, post-TPO training, the overall image attention weights for response tokens significantly increase, especially for visual-anchored tokens (e.g., table, cord). This underscores our method’s efficacy in improving the model’s capacity to integrate visual information during response generation, thus mitigating the hallucination issue.
Self-Calibration : To illustrate that our method enables the model to progressively enhance its focus on visual information through continuous self-calibration during training, we present the evolution of scores for positive and negative samples, as calculated by Equation (6), across various training steps. With $a=0.5$, it follows that $c_{y_{i}}\in(0.5,1.5)$. As shown in Figure 5, the scores for positive samples gradually approach their maximum values, while those for negative samples approach their minimum values, indicating convergence. This trend illustrates the self-calibrating effect of our method, which ultimately enhances the model’s ability to focus on visual information.
Figure 5: The curve of changes in self-calibrated rewards for positive and negative samples over training steps, with a sample point taken every 10 steps.
5 Conclusion
In this study, we propose TPO, a novel token preference optimization framework, aimed at mitigating hallucinations in LVLMs. TPO incorporates a self-calibrated visual-anchored reward mechanism that automatically identifies "vision-anchored tokens" and adaptively assigns appropriate rewards to them. By adding noise to the visual input and capturing changes in the generation probability of each token, TPO computes a score indicating each token’s relevance to visual information. Subsequently, a self-calibration process adjusts these scores to ensure that: 1) Rewards for positive examples exceed those for negative examples. 2) The final rewards enhance the focus on visual information for all of "vision-anchored tokens" in preference data. Based on the self-calibrated visual-anchored reward, TPO can perform more efficient token-level preference alignment optimization for LVLMs. Experimental results have proved that TPO not only alleviates the hallucination problem but also strengthens the model’s attention to visual input when generating responses.
6 Limitation
Although our method has achieved outstanding performance in addressing the hallucination problem, the self-calibrated visual-anchored rewards approach we used in this paper can be extended to even broader areas. By altering the way noise is added to images, we can shift from adding noise to the entire image to adding noise to specific key objects. It can enable the model to specifically improve its focus on image information in certain domains, thus having extensive industrial applications. We will continue to expand in this direction, and we believe that the technology we have proposed in this paper has a vast space for further development and application.
7 Ethic Statement
The main purpose of this article is to alleviate the hallucination problem in LVLM using reinforcement learning method. By employing a self-calibrated visual-anchored reward approach, we propose the TPO method, which significantly addresses the hallucination issue and helps the model connect with more visual information. All the models and datasets we used are open source, so we believe that the work in this paper does not pose any potential threats.
Appendix A Setup Details
In our experiments, we trained the LLaVA-v1.5 model. For our TPO method and the vanilla DPO method, we set the maximum learning rate to 5e-8 on the 7B version and trained for 4 epochs. We set the maximum learning rate to 2e-7 on the 7B version and trained for 4 epochs. The RLHF-V training was set according to the paper Yu et al. (2024b). All parts requiring GPT-4 evaluation use the GPT-4-0613 version, and the MM-Vet testing is conducted on the official evaluation website.
Our experiments were all conducted on a server equipped with 8 GPUs; in specific cases (such as the 13B model), we utilized 32 GPUs. For the hyperparameter setting, all hyperparameters are consistent with those of our main experiment. Moreover, the level of diffusion noise in our model is represented by a formula $\xi=Sigmoid(l_{t})\times(0.5\times 10^{-2}-10^{-5})+10^{-5}$, where $l_{t}$ is a list of 1,000 numbers taken at equal intervals over the interval $[-6,6]$.
The cases in Figure 1 and Figure 4 come from benchmarks Sun et al. (2023), while the cases in Figure 2 come from the RLHFV training set Yu et al. (2024b).
Appendix B Experimental Results
In this section, we will provide a detailed presentation of further ablation studies on the noise steps and $a$ in Eq 6, as shown in Table 4 and Table 5. The optimal performance for noise steps is attained at the 500-steps. This level of medium corruption allows the model to capture the general outline of the image while neglecting finer details, which can lead to the generation of hallucinations in the visual-anchored tokens. and The Optimal performance for $a$ is $0.5$, when the visual-anchored score $s=0$, we get $c=1$, not introducing additional reward signals can yield better outcomes.

### v3

@@
 Extensive experimental results have manifested the state-of-the-art performance of the proposed TPO.
-For example, by building on top of LLAVA-1.5-7B, our TPO boosts the performance absolute improvement for hallucination benchmarks.
+For example, by building on top of LLaVA and Qwen, our TPO boosts the performance absolute improvement for hallucination benchmarks.
 Token Preference Optimization with Self-Calibrated Visual-Anchored Rewards for Hallucination Mitigation
@@
   • Extensive experiments on the popular hallucination benchmarks demonstrate the state-of-the-art performance of the proposed TPO.
-Methods Visual- Anchored Token- level Fine-grained Annotations
+Methods Visual- Anchored Token- level Non Fine-grained Annotations
 DPO
@@
 CSR
+V-DPO
 RLHF-V
-V-DPO
 TPO (Ours)
@@
 (2023); Lee et al.
-(2023); Xia et al.
 (2023); Shao et al.
@@
 While these methods can lead to more accurate responses, they do not fundamentally resolve the issue of inadequate visual information association in LVLMs.
+Figure 2: Outline of our TPO pipline.
+The process is divided into three parts for each data at every training step.
+First, 1) add noise to the image, then, 2) calculate Self-Calibrated Visual-Anchored Rewards, and finally 3) perform Token Preference Optimization.
+At the end of each training step, we calibrate the model and calculate new Visual-Anchored Rewards for the next step.
 2.2 Preference Learning Methods
@@
 3 Methodology
-Figure 2: Outline of our TPO pipline.
-The process is divided into three parts for each data at every training step.
-First, 1) add noise to the image, then, 2) calculate Self-Calibrated Visual-Anchored Rewards, and finally 3) perform Token Preference Optimization.
-At the end of each training step, we calibrate the model and calculate new Visual-Anchored Rewards for the next step.
-3.1 Motivation
-As illustrated in Figure 1, the generation of each token requires differing levels of integration with visual information.
-Tokens that are closely correlated with visual inputs are designated as "visual-anchored tokens" in this paper.
-It is observed that the generation of such visual-anchored tokens can be more susceptible to hallucinations.
-Furthermore, directly assigning rewards to entire generated sequences can lead to tokens independent of visual information consuming undue optimization resources.
-This misallocation can reduce the supervisory focus on visual-anchored tokens.
-To enhance training efficiency and ensure the model attends to relevant visual features, it is crucial to assign token-level rewards based on the reliance on visual information during training.
-To achieve these goals, the proposed TPO mitigates hallucination from the following three key aspects: 1) scalable token-level rewards to optimize alignment; 2) automatic identification of visual-anchored tokens by the model itself; 3) self-calibration of reward signals based on visual correlation.
+The schematic illustration of the proposed TPO is demonstrated in Figure 2.
+In Sec.
+3.1, we present the preliminaries including the definition and off-policy optimization of DPO.
+Then we detail the visual-anchored rewards and token preference optimization loss in Sec.
+3.2 and Sec.
+3.3, respectively.
 3.1 Preliminaries
@@
 (2024) is designed to directly maximize the reward margin between positive and negative responses to align human preferences.
-Specifically, given a textual input $x$, a visual input $v$, a negative response $y_{l}$, and a preferred positive response $y_{w}$, the reward function is defined as $r(x,v,y_{l}/y_{w})$.
-According to the Bradley-Terry Bradley and Terry (1952) model, the probability of each response is as follows:
-$p(y_{w}>y_{l})=\sigma(r(x,v,y_{w})-r(x,v,y_{l}))$ (1)
-where the sigmoid function is denoted as $\sigma(\cdot)$.
-In this paper, we focus on the off-policy DPO method, where the probability of preference data is derived from the optimal policy rather than the reward model. The reward function is designed as:
-$r(x,v,y)=\beta\log\frac{\pi_{\theta}(y|x,v)}{\pi_{\text{ref}}(y|x,v)}$ (2)
-In Equation (2), assuming a response $y\in\{y_{w},y_{l}\}$ has $N$ tokens $(y_{1},...,y_{i},...,y_{N})$, the output distribution of the current policy model $\pi_{\theta}(y|x,v)$ and the reference model $\pi_{\text{ref}}{(y|x,v)}$ can be expressed as $\prod_{y_{i}}p(y_{i}|x,v)$.
+Given a textual input $x$, a visual input $v$, a negative response $y_{l}$, and a preferred positive response $y_{w}$, the reward function $r(x,v,y_{l}/y_{w})$ is defined as follows.
+$r(x,v,y)=\beta\log\frac{\pi_{\theta}(y|x,v)}{\pi_{\text{ref}}(y|x,v)},$ (1)
+where $\pi_{\text{ref}}{(y|x,v)}$ and $\pi_{\theta}(y|x,v)$ respectively represent the reference model and current policy model.
 On this basis, the formulation of a maximum likelihood objective is defined as:
-$\mathcal{L}_{DPO}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,v,y_{w},y_{l})\sim D}\Bigg{[}\log\sigma$ (3)
-$\Bigg{(}\beta\log\frac{\pi_{\theta}(y_{w}|x,v)}{\pi_{\text{ref}}(y_{w}|x,v)}-\beta\log\frac{\pi_{\theta}(y_{l}|x,v)}{\pi_{\text{ref}}(y_{l}|x,v)}\Bigg{)}\Bigg{]}$
-It is demonstrated that DPO assigns equal rewards to all tokens in the responses.
-When applied to LVLMs for mitigating hallucinations, it lacks emphasis on visual-anchored tokens.
-Therefore, we propose a novel approach named TPO to mainly solve this problem.
-3.3 TPO
-3.3.1 Visual-Anchored Rewards
-We propose a self-calibrated visual-anchored reward generation mechanism based on Equation (2).
-On the one hand, it can automatically identify visual-anchored tokens in the response during training.
-On the other hand, it adaptively and iteratively assigns different rewards for tokens according to their visual reliance.
-The proposed rewards can not only strengthen the attention to visual information for token generation but also align preferences.
-To construct the visual-anchored rewards, we first introduce a score $s$ to measure the visual reliance degree of each token.
-For this purpose, we add noise into the embedding of the input image $v$ to obtain the corrupted image $v_{c}$:
-$v_{c}(k)=\sqrt{\bar{\xi}_{k}}\cdot v+\sqrt{1-\bar{\xi}_{k}}\cdot\epsilon$ (4)
-Here, we mainly refer Zhou et al.
-(2024a) and set the noise steps $k=500$, $\bar{\xi}_{k}=\prod_{i=0}^{k}\xi_{i}$, $\xi$ is a predefined noise parameter derived from a list with 1,000 equally spaced elements.
-[^1] More details can be found in Appendix A.
-Subsequently, we capture the changes in the probability distribution of each generated token before and after adding noise.
-For token $y_{i}$ of the response $y$, the calculation of $s_{y_{i}}$ is formulated as Equation (5).
-Notably, here $p_{log}$ refers to the raw logits output of the model, before applying the softmax transformation.
-$s_{y_{i}}=p_{log}(y_{i}|x,v,y_{<i})-p_{log}(y_{i}|x,v_{c},y_{<i})$ (5)
-We have shown an example of the distribution of $s$ before and after training in Figure 1, it is shown that visual-anchored tokens will obtain higher scores.
-In other words, $s$ has the ability to reflect the visual relevance of each token $y_{i}$, this also lays the foundation for the ability of our visual-anchored rewards to self-calibrate based on the visual correlation of tokens.
-Furthermore, we introduce a self-calibration process to generate the final visual-anchored rewards.
+$\mathcal{L}_{DPO}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,v,y_{w},y_{l})\sim D}\Big{[}\log\sigma$ (2)
+$\Big{(}\beta\log\frac{\pi_{\theta}(y_{w}|x,v)}{\pi_{\text{ref}}(y_{w}|x,v)}-\beta\log\frac{\pi_{\theta}(y_{l}|x,v)}{\pi_{\text{ref}}(y_{l}|x,v)}\Big{)}\Big{]},$
+where $\sigma(\cdot)$ denotes the sigmoid function.
+3.2 Visual-Anchored Rewards
+Different to the equal confidence for each token in DPO, we propose a visual-anchored by measuring the token-wise visual reliance.
+Specifically, we firstly add noise into the embedding of the input image $v$ in a total $k$ steps to obtain the corrupted image $v_{c}$:
+$v_{c}(k)=\sqrt{\bar{\xi}_{k}}\cdot v+\sqrt{1-\bar{\xi}_{k}}\cdot\epsilon,$ (3)
+where $\xi$ is a predefined noise parameter derived from a list with 1,000 equally spaced elements [^1] More details can be found in Appendix A, and experimental analysis can be found in Appendix E.
+$\bar{\xi}_{k}$ is a cumulant, i.e., $\bar{\xi}_{k}=\prod_{i=0}^{k}\xi_{i}$.
+Subsequently, the difference of generated token distribution is computed as follow:
+$s_{y_{i}}=p_{log}(y_{i}|x,v,y_{<i})-p_{log}(y_{i}|x,v_{c},y_{<i}),$ (4)
+where $s_{y_{i}}$ denotes the distribution difference of the token $y_{i}$ of the response $y$.
+$p_{log}$ refers to the raw logits output of the model, before applying softmax normalization.
+One example case is demonstrated in Figure 1, which demonstrates that $s$ reflects the visual relevance of each token $y_{i}$.
+Then, a self-calibration process is proposed to generate the final visual-anchored rewards $c_{y_{i}}$.
+$c_{y_{i}}=\begin{cases}a+\sigma(s_{y_{i}})&\text{if }y_{i}\in y_{w}\\ a+1-\sigma(s_{y_{i}})&\text{if }y_{i}\in y_{l}\end{cases}$ (5)
+where $a$ is a margin value.
+We set $a=0.5$ in Equation (5), so that when $s=0$, $c=1$, the rewards will not take effect.
 This process aims to ensure that positive samples receive higher rewards than negative samples while optimizing the visual relevance of visual-anchored tokens in all responses.
-Formally, the process is defined as follows, where the sigmoid function $\text{sigmoid}(\cdot)$ is used for normalization.
-$c_{y_{i}}=\begin{cases}a+\text{sigmoid}(s_{y_{i}})&\text{if }y_{i}\in y_{w}\\ a+1-\text{sigmoid}(s_{y_{i}})&\text{if }y_{i}\in y_{l}\end{cases}$ (6)
-The calculated $c_{y_{i}}$ is adopted as the final reward in TPO, where $c_{y_{i}}\in(a,a+1)$.
-In this way, $c$ and $s$ are positively correlated for positive examples, whereas for negative ones, the opposite holds true.
-In particular, $p(y_{i}|x,v,y_{<i})=p(y_{i}|x,v_{c},y_{<i})$, means that the token does not associate any visual information.
-In this case, no visual-anchored reward should be assigned for this token.
-To this end, we set $a=0.5$ in Equation (6), so that when $s=0$, $c=1$, the rewards will not take effect.
-3.3.2 Token Preference Optimization
+3.3 Token Preference Optimization
 After obtaining the reward $c_{y_{i}}$ to $y_{i}$, the output cumulative distribution can be calculated:
-$\pi(y|x,v)=\prod_{y_{i}\in\mathcal{Y}}c_{y_{i}}$ (7)
+$\pi^{v}(y|x,v)=\prod_{y_{i}\in\mathcal{Y}}c_{y_{i}}$ (6)
 Especially, when $c_{y_{i}}=1$, the probability of $y_{i}$ will not be accumulated.
 By multiplying the probability distribution with the visual-anchored rewards, we obtain a novel KL-constrained reward maximization objective:
-$\max_{\pi}E_{(x,v,y)}\Bigg{[}\,r^{\prime}(x,v,y)-\beta D_{KL}\Bigg{(}\left[\pi_{\theta}(y|x,v)\right.$ (8)
-×$\left.\times\pi_{\theta}^{v}(y|x,v)\right]\Bigg{\|}\left[\pi_{\text{ref}}(y|x,v)\times\pi_{\text{ref}}^{v}(y|x,v)\right]\Bigg{)}\Bigg{]}$
+$\max_{\pi}E_{(x,v,y)}\Big{[}\,r^{\prime}(x,v,y)-\beta D_{KL}\Big{(}\pi_{\theta}(y|x,v)$ (7)
+$\cdot\pi_{\theta}^{v}(y|x,v),\pi_{\text{ref}}(y|x,v)\cdot\pi_{\text{ref}}^{v}(y|x,v)\Big{)}\Big{]},$
+where $D_{KL}(\cdot,\cdot)$ denotes the KL divergency computation.
+$\pi_{\theta}^{v}(y|x,v)$ and $\pi_{\text{ref}}^{v}(y|x,v)$ are calculated using the policy model and the reference model, respectively.
 Thus, the optimal solution formula for the maximization objective of the KL-constrained reward is as follows:
-$\pi_{\theta}(y|x,v)\times\pi_{\theta}^{v}(y|x,v)=\frac{1}{Z(x)}\left[\pi_{\text{ref}}(y|x,v)\times\right.$ (9)
-$\left.\pi_{\text{ref}}^{v}(y|x,v)\right]\exp\left(\frac{1}{\beta}r^{\prime}(x,v,y)\right)$
-Rearranging the Eq (9), we obtain the new reward function:
+$\pi_{\theta}(y|x,v)\cdot\pi_{\theta}^{v}(y|x,v)=\frac{1}{Z(x,v)}\pi_{\text{ref}}(y|x,v)\cdot$ (8)
+$\pi_{\text{ref}}^{v}(y|x,v)\exp\big{(}\frac{1}{\beta}r^{\prime}(x,v,y)\big{)}.$
+The partition function of Eq (8) is as follows.
+$Z(x,v)=\sum_{y}$ $\pi_{\text{ref}}(y|v,x)\cdot\pi_{\text{ref}}^{v}(y|x,v)$ (9)
+$\cdot\exp\big{(}\frac{1}{\beta}r^{\prime}(x,v,y)\big{)}$
+Rearranging Eq (8), we obtain the reward function:
 $r^{\prime}(x,v,y)$ (10)
-$=\beta\log\frac{\pi_{\theta}(y|x,v)\times\pi_{\theta}^{v}(y|x,v)}{\pi_{\text{ref}}(y|x,v)\times\pi_{\text{ref}}^{v}(y|x,v)}+\beta Z(x)$
-$=\beta\sum_{y_{i}\in\mathcal{Y}}\left[\log\left(p_{\theta}(y_{i}|x,v,y_{<i})\cdot c_{\theta}\right)\right.$
-$\quad\left.-\log\left(p_{\text{ref}}(y_{i}|x,v,y_{<i})\cdot c_{\text{ref}}\right)\right]+\beta Z(x)$
-$=\beta\sum_{y_{i}\in\mathcal{Y}}\left[\log p_{\theta}(y_{i}|x,v,y_{<i})-\right.$
-$\quad\left.\log p_{\text{ref}}(y_{i}|x,v,y_{<i})+\log\frac{c_{\theta}}{c_{\text{ref}}}\right]+\beta Z(x)$
-Compared to the original reward function in DPO (Equation (2)), we multiply each $p(y_{i}|x,v,y_{<i})$ by the generated visual-anchored rewards $c_{y_{i}}$ at the token level.
-For $c_{\theta}$, its computation is continuously updated at each step during training as the model changes.
-In the calculation of each token in the entire reward function, we add a term $log\frac{c_{\theta}}{c_{\text{ref}}}\in(-log3,log3)$, which has a reasonable upper and lower bound.
+$=\beta\log\frac{\pi_{\theta}(y|x,v)\cdot\pi_{\theta}^{v}(y|x,v)}{\pi_{\text{ref}}(y|x,v)\cdot\pi_{\text{ref}}^{v}(y|x,v)}+\beta Z(x,v)$
+$=\beta\sum_{y_{i}\in y}\left[\log\left(p_{\theta}(y_{i}|x,v,y_{<i})\cdot{\color[rgb]{1,0,0}c_{y_{i}}^{\theta}}\right)\right.$
+$\quad\left.-\log\left(p_{\text{ref}}(y_{i}|x,v,y_{<i})\cdot{\color[rgb]{0,0,1}c_{y_{i}}^{\text{ref}}}\right)\right]+\beta Z(x,v)$
+$=\beta\sum_{y_{i}\in y}\Big{[}\log p_{\theta}(y_{i}|x,v,y_{<i})-\log p_{\text{ref}}(y_{i}|x,v,y_{<i})$
+$\quad+\log\frac{{\color[rgb]{1,0,0}c_{y_{i}}^{\theta}}}{{\color[rgb]{0,0,1}c_{y_{i}}^{\text{ref}}}}\Big{]}+\beta Z(x,v),$
+where $c_{y_{i}}^{\theta}$ and $c_{y_{i}}^{\text{ref}}$ represent the token reward calculated using the policy model and the reference model, respectively.
+Compared to the original reward function in DPO (Eq (1)), we multiply each $p(y_{i}|x,v,y_{<i})$ by the generated visual-anchored rewards $c_{y_{i}}$ at the token level.
+$c_{y_{i}}^{\theta}$ is continuously updated at each step during training as the model changes.
+To calculate each token in the entire reward function, we add a term $log\frac{c_{y_{i}}^{\theta}}{c_{y_{i}}^{\text{ref}}}\in(-log3,log3)$, which has a reasonable upper and lower bound.
 For positive samples, this term is expected to increase, while for negative samples, it is expected to decrease.
-Due to the different methods of calculating $c_{y_{i}}$ that we set in Equation (6), this will encourage the increase of $s_{y_{i}}$ during the training process, making the token generation focus more on visual information.
-Thus, following the Bradley-Terry model, when given the positive and negative samples $\mathcal{D}=\{x^{(i)},v^{(i)},y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{N}$, we obtain our maximum likelihood objective:
-$\mathcal{L}_{TPO}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,v,y_{w},y_{l})\sim D}\bigg{[}\log\sigma$ (11)
-$\quad\Bigg{(}\beta\log\frac{\pi_{\theta}(y_{w}\big{|}x,v)\times\pi_{\theta}^{v}(y_{w}\big{|}x,v)}{\pi_{\text{ref}}(y_{w}\big{|}x,v)\times\pi_{\text{ref}}^{v}(y_{w}\big{|}x,v)}-$
-$\quad\beta\log\frac{\pi_{\theta}(y_{l}\big{|}x,v)\times\pi_{\theta}^{v}(y_{l}\big{|}x,v)}{\pi_{\text{ref}}(y_{l}\big{|}x,v)\times\pi_{\text{ref}}^{v}(y_{l}\big{|}x,v)}\Bigg{)}\bigg{]}$
+Due to the different methods of calculating $c_{y_{i}}$ that we set in Eq (5), this will encourage the increase of $s_{y_{i}}$ during the training process, making the token generation focus more on visual information.
+Thus, following the Bradley-Terry model, when given the positive and negative samples $\mathcal{D}=\{x^{(k)},v^{(k)},y_{w}^{(k)},y_{l}^{(k)}\}_{k=1}^{N}$, we obtain our maximum likelihood objective:
+$\mathcal{L}_{TPO}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,v,y_{w},y_{l})\sim D}\Big{[}\log\sigma$ (11)
+$\quad\Big{(}\beta\log\frac{\pi_{\theta}(y_{w}\big{|}x,v)\cdot\pi_{\theta}^{v}(y_{w}\big{|}x,v)}{\pi_{\text{ref}}(y_{w}\big{|}x,v)\cdot\pi_{\text{ref}}^{v}(y_{w}\big{|}x,v)}-$
+$\quad\beta\log\frac{\pi_{\theta}(y_{l}\big{|}x,v)\cdot\pi_{\theta}^{v}(y_{l}\big{|}x,v)}{\pi_{\text{ref}}(y_{l}\big{|}x,v)\cdot\pi_{\text{ref}}^{v}(y_{l}\big{|}x,v)}\Big{)}\Big{]}$
+$=\mathcal{L}_{DPO}(\pi_{\theta};\pi_{\text{ref}})+\mathbb{E}_{(x,v,y_{w},y_{l})\sim D}\Big{[}\log\sigma$
+$\quad\Big{(}\beta\log\frac{\pi_{\theta}^{v}(y_{w}\big{|}x,v)}{\pi_{\text{ref}}^{v}(y_{w}\big{|}x,v)}-\beta\log\frac{\pi_{\theta}^{v}(y_{l}\big{|}x,v)}{\pi_{\text{ref}}^{v}(y_{l}\big{|}x,v)}\Big{)}\Big{]}$
+According to Eq (10), we can deduce as follows.
+$\mathcal{L}_{TPO}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,v,y_{w},y_{l})\sim D}\Big{[}\log\sigma$ (12)
+$\quad\Big{(}\beta\sum_{y_{w_{i}}\in y_{w}}\Big{[}\log\left(p_{\theta}(y_{w_{i}}|x,v,y_{w_{<i}})\right)$
+$\quad-\log p_{\text{ref}}(y_{w_{i}}|x,v,y_{w_{<i}})+\log\frac{c_{y_{w_{i}}}^{\theta}}{c_{y_{w_{i}}}^{\text{ref}}}\Big{]}$
+$+\sum_{y_{l_{i}}\in y_{l}}\Big{[}\log\left(p_{\theta}(y_{l_{i}}|x,v,y_{l_{<i}})\right)$
+$\quad-\log p_{\text{ref}}(y_{l_{i}}|x,v,y_{l_{<i}})+\log\frac{c_{y_{l_{i}}}^{\theta}}{c_{y_{l_{i}}}^{\text{ref}}}\Big{]}\Big{)}\Big{]}$
+where $c_{y_{w_{i}}}^{\theta}$ and $c_{y_{w_{i}}}^{\text{ref}}$ represent the token reward calculated for $y_{w}$ using the policy model and the reference model, respectively.
+The same applies to $c_{y_{w_{i}}}^{\theta}$, $c_{y_{w_{i}}}^{\text{ref}}$ and $y_{l}$.
 Method AMBER MMHal HallusionBench General Benchmarks
-Acc F1 Score Hall $\downarrow$ Easy Hard aAcc SEED MMB LLaVA MM-Vet
+Acc F1 Score Hal $\downarrow$ Easy Hard aAcc SEED MMB LLaVA MM-Vet
 LLaVA-1.5-7B 71.7 74.3 2.01 61.46 42.64 41.16 47.21 66.1 73.3 65.6 31.6
@@
 Table 2: Performence of LLaVA-1.5 on hallucination and general benchmarks.
-Score and Hall refer to the overall GPT-4 score and hallucination rate, respectively.
+Score and Hall refer to the overall GPT-4 Achiam et al.
+(2023) score and hallucination rate, respectively.
 Easy represents the accuracy of with original images, hard represents the accuracy with manually edited challenging images, and aAcc is the average accuracy for each question.
@@
 (2024b) are based on our testing of their open-source model weights, while the results for V-DPO Xie et al.
-(2024) are taken from previous work.
-(bold: the best score; underline: the second best)
+(2024) are taken from previous work (bold: the best score; underline: the second best)
+Method AMBER MMHal HallusionBench
+[…]
+Qwen2-vl 86.5 90.0 3.5 29.0 67.0 48.8 64.0
++DPO 86.5 90.0 3.7 28.1 67.3 49.3 64.5
++TPO 86.4 89.9 4.2 18.8 67.9 50.0 65.2
+Table 3: Performence of Qwen2-vl-7B on hallucination benchmarks.
 4 Experiment
 4.1 Setup
-Aligning with previous DPO-based approaches on hallucination mitigation, we adopt the popular LVLM, LLaVA-1.5 Liu et al.
+Aligning with previous DPO-based approaches on hallucination mitigation, we mainly adopt the popular LVLM, LLaVA-1.5 Liu et al.
 (2024b), as the backbone model to validate the effectiveness of our TPO.
-As for the dataset, we directly utilize the preference pairs provided by RLHF-V (5k) without their fine-grained human annotations.
-Benchmarks: We primarily conduct the experiments on three hallucination benchmarks: (1) AMBER Wang et al.
-(2023): a multi-dimensional hallucination benchmark with over 15k samples.
-We mainly focus on its discriminative part and report Accuracy and F1 metrics referencing Yu et al.
-(2024c); (2) MMHal-Bench Sun et al.
-(2023): it measures the hallucination rate and informativeness of responses.
-(3) HallusionBench Guan et al.
-(2024): it evaluates visual illusions and knowledge hallucinations through systematically structured discriminative tasks.
-We also test the performance on four general evaluation benchmarks: (1) SEED Bench Li et al.
-(2023a): a benchmark for LVLMs on generative comprehension.
-(2) MMBench Liu et al.
-(2025): a comprehensive benchmark designed to evaluate the capabilities across various tasks and modalities.
-(3) LLaVA Bench Liu et al.
-(2024c): a benchmark for evaluating multi-modal conversation, detailed description, and complex reasoning.
-(4) MM-Vet Yu et al.
-(2023): a benchmark assessing for integrated capabilities.
-Baselines: In this work, we mainly compare TPO with the SFT model of LLaVA-1.5-7B, as well as with DPO and V-DPO Xie et al.
+Furthermore, to evaluate the effectiveness of TPO on more advanced and powerful model, we implement TPO training based on Qwen2-vl Wang et al.
+(2024), and compare it with the DPO method.
+For the dataset, we directly utilize the preference pairs provided by RLHF-V (5K) without their fine-grained human annotations.
+Benchmarks
+We primarily conduct the experiments on three hallucination benchmarks: AMBER Wang et al.
+(2023), MMHal-Bench Sun et al.
+(2023), and HallusionBench Guan et al.
+(2024).
+In this section, we mainly focus on AMBER’s discriminative task and report the accuracy and F1 metrics referencing Yu et al.
+(2024c).
+In addition, we provide the results of its Chair metric in Appendix D.
+Moreover, we also evaluate the performance of TPO on four general benchmarks: SEED Bench Li et al.
+(2023a), MMBench Liu et al.
+(2025), LLaVA Bench Liu et al.
+(2024c) and MM-Vet Yu et al.
+(2023).
+These benchmarks are used to evaluate the performance of the models on general tasks after hallucination alignment.
+Baselines
+We mainly compare TPO with the LLaVA-1.5-7B SFT model, as well as with the DPO and V-DPO Xie et al.
 (2024) methods trained using RLHF-V Yu et al.
@@
 (2024a).
-Moreover, to evaluate the effectiveness and robustness of TPO as the model size increases, we further compare it with the LLaVA-1.5-13B SFT model, and also against DPO and RLHF-V methods.
-Figure 3: Performance curves with the change of the noise steps-(a) and the change of parameter $a$-(b), We separately present the F1 of AMBER, the hallucination rate of MMHAL, the aACC of HallusionBench, and the average value of the general benchmarks.
-More detailed metrics can be found in the Appendix B.
+Moreover, to evaluate the effectiveness and robustness of TPO as the model size increases, we further evaluate the performance of TPO on the LLaVA-1.5-13B model and compared it with DPO.
+Additionally, to demonstrate the advantages of TPO, we reproduced the strong baseline method, RLHF-V, on LLaVA-1.5-13B and conducted a comparison.
+To further demonstrate the effectiveness of TPO on more advanced and powerful model, we additionally employ Qwen2-vl-7B as the baseline mode and compare our TPO with DPO.
+Method AMBER MMHal HallusionBench General Benchmarks
+Acc F1 Score Hal $\downarrow$ Easy Hard aAcc SEED MMB LLaVA MM-Vet
+LLaVA-1.5-7B 71.70 74.3 2.01 61.46 42.64 41.16 47.21 66.1 73.3 65.6 31.6
+Only Win 79.10 84.5 2.24 56.25 44.62 46.05 50.40 66.6 73.6 69.8 31.7
+Only Loss 79.20 84.8 2.33 53.13 42.20 47.91 49.87 66.6 73.5 70.7 32.0
+Opposite 75.30 80.7 1.91 64.58 42.42 45.58 48.63 65.6 73.1 68.9 32.1
+TPO (Ours) 79.30 85.0 2.47 51.04 41.76 48.37 50.22 66.6 73.6 70.2 33.0
+Table 4: Ablation Studies. Performence of LLaVA-1.5 on hallucination and general benchmarks.
+Average score Noun/Adj Others
+Ground Truth 1.83 0.90
+Ground Truth (TPO) 5.72 4.87
+Response of LLaVA 1.48 0.83
+Response of LLaVA+TPO (TPO) 5.67 4.59
+Table 5: Average score from Equation 4 of Noun/Adj tokens and other tokens.
+Here, Ground Truth and Ground Truth (TPO) represent the scores calculated for the ground truth answer using LLaVA-1.5-7B and LLaVA-1.5-7B+TPO.
+Response of LLaVA and LLaVA+TPO (TPO) correspond to the outputs before and after TPO training and the scores calculated by LLaVA-1.5-7B and LLaVA-1.5-7B+TPO, corresponding to Fiure 1.
 4.2 Main Results
@@
 This indicates that, compared to other methods, our approach enables the model to focus more on visual information rather than the textual prior knowledge to provide accurate answers.
-On general benchmarks, compared to other methods, our approach remains stable against the original LLaVA model and achieves the greatest improvement on most benchmarks.
+On general benchmarks, our approach remains stable against the original LLaVA model and achieves the greatest improvement on most benchmarks.
 We attribute it to that our method helps the model associate with more visual information when answering questions.
 This shows that our approach can improve hallucination issues while maintaining good performance on general evaluation tasks.
-Method AMBER MMHal HallusionBench General Benchmarks
-Acc F1 Score Hall $\downarrow$ Easy Hard aAcc SEED MMB LLaVA MM-Vet
-LLAVA-1.5-7B 71.70 74.3 2.01 61.46 42.64 41.16 47.21 66.1 73.3 65.6 31.6
-Only Win 79.10 84.5 2.24 56.25 44.62 46.05 50.40 66.6 73.6 69.8 31.7
-Only Loss 79.20 84.8 2.33 53.13 42.20 47.91 49.87 66.6 73.5 70.7 32.0
-Opposite 75.30 80.7 1.91 64.58 42.42 45.58 48.63 65.6 73.1 68.9 32.1
-TPO (Ours) 79.30 85.0 2.47 51.04 41.76 48.37 50.22 66.6 73.6 70.2 33.0
-Table 3: Ablation Studies. Performence of LLaVA-1.5 on hallucination and general benchmarks. (bold: the best score)
-Figure 4: Comparison of attention weights for LLaVA-1.5 before and after TPO training.
-Each horizontal line represents the mean of that data set.
-The blue section on the left response incorrectly, with many ’visual-anchored tokens’ tokens having high attention weights but resulting in hallucinated responses (e.g.
-USB, cloud).
+The results also indicate that TPO introduces performance fluctuations on the LLaVABench and MMVet metrics for LLaVA-1.5-13B.
+This occurs because the training data primarily focuses on aligning preferences within the context of hallucinations, which inevitably leads to slight overfitting.
+Although augmenting the dataset with task-specific data could mitigate this issue, it is not the research focus of this paper.
+4.3 Results on Qwen2-vl
+As Table 3 shown, we report the results on the key metrics of three hallucination benchmarks.
+The results indicate that our TPO outperforms DPO on most benchmarks.
+On Qwen2-vl, which has strong inherent capabilities, using 5K RLHF-V data for DPO alignment barely improves the performance.
+However, introducing TPO leads to a significant further enhancement.
+This not only demonstrates the effectiveness of TPO on other powerful models but also shows that TPO’s token-level visual anchor rewards brings higher data utilization efficiency.
+Although Qwen2-vl shows almost no room for improvement on AMBER, TPO still achieves significant improvement on the Chair metric, as shown on Figure 5.
+TPO can capture and learn more subtle preferences from the data.
+Figure 3: Comparison of attention weights for LLaVA before and after TPO training.
+Each horizontal line represents the mean of that data.
+The blue section response incorrectly, with many ’visual-anchored tokens’ tokens having high attention weights but resulting in hallucinated responses (e.g.
+USB).
 The red section on the right answered correctly.
+Figure 4: The curve of changes in self-calibrated rewards for positive and negative samples over training steps, with a sample point taken every 10 steps.
 4.4 Ablation Studies
-To further demonstrate how our method helps improve model performance by reducing hallucinations, we conduct ablation experiments for the noisy steps, the introduced parameter $a$, and the self-calibrated process for rewards.
-These experiments are performed on LLaVA-1.5-7B consistently with the main experiments, including hallucination and general evaluation benchmarks.
-Noise Steps : We ablate on the noise steps in Figure 3 (a).
-As shown, the optimal performance is achieved at the step of 500.
-This medium corruption enables the model to grasp the general outline of the image while missing the detailed contents, which is prone to generate hallucinations of the visual-anchored tokens.
-Parameter $a$ : We conduct experiments by varying the parameter $a$ introduced in Equation (6) with the results shown in Figure 3 (b).
-By setting $a=[0,0.5,1]$, we observed consistently good performance across all configurations.
-This suggests that effective performance is achieved as long as the reward mechanism successfully highlights token differences and identifies visually anchored tokens.
-Notably, the best overall results are obtained with $a=0.5$, validating our proposed method and hypothesis.
-This indicates that when the visual-anchored score $s=0$, setting $c=1$, not introducing additional reward signals can yield better outcomes.
-Visual-Anchored Rewards : Furthermore, Table 3 demonstrates that TPO can enhance model performance when rewards are assigned separately to positive and negative samples, achieving results comparable to those obtained by rewarding both simultaneously.
+Visual-Anchored Rewards
+Table 4 demonstrates that TPO can enhance model performance when rewards are assigned separately to positive and negative samples, achieving results comparable to those obtained by rewarding both simultaneously.
 However, by providing opposite rewards to positive and negative samples, where rewards are negatively correlated with the visual relevance of positive samples and positively correlated with that of negative samples, TPO’s performance significantly deteriorates.
@@
 This further underscores the validity of the designation of visual-anchored rewards.
-By integrating these selection methods, we achieve superior overall performance, thereby highlighting the value and efficacy of our approach.
-This integration underscores both the simplicity and effectiveness of our method in achieving robust results.
+Hyperparameters
+To optimize the hyperparameters in TPO, we perform comparative experiments on the noise steps (Section 3.2) and parameter $a$ in Equation 5.
+As the Figure 5 shown, the model performs best with 500 noise steps.
+Testing $a$ across the range $a=[0,0.5,1]$, we find that $a$=0.5 produces the best outcome.
+This supports our hypothesis that setting $s=0$ and $c=1$, without adding extra reward signals, leads to superior performance [^2] The more detail results and discussions are provided in Appendix..
+Figure 5: Performance curves with the change of the noise steps-(a) and the change of parameter $a$-(b), We separately present the F1 of AMBER, the hallucination rate of MMHAL, the aACC of HallusionBench, and the average value of the general benchmarks.
+More detailed metrics can be found in the Appendix B.
 4.5 Analysis
-Attentions : To further substantiate the effectiveness of our method in enhancing visual alignment, we visualize the relevance between response and image tokens, both before and after training with TPO.
-In this statistic, the relevance is quantified using the sum of attention weights.
-As depicted in Figure 4, post-TPO training, the overall image attention weights for response tokens significantly increase, especially for visual-anchored tokens (e.g., table, cord).
-This underscores our method’s efficacy in improving the model’s capacity to integrate visual information during response generation, thus mitigating the hallucination issue.
-Self-Calibration : To illustrate that our method enables the model to progressively enhance its focus on visual information through continuous self-calibration during training, we present the evolution of scores for positive and negative samples, as calculated by Equation (6), across various training steps.
+Visual-Anchored Rewards
+As Figure 1 shown, the proposed visual-anchored rewards can reflect the degree to which a token depends on visual information.
+To further prove this statement, we construct the analysis experiment on the MMhal dataset as shown in Table 5.
+Intuitively, nouns and adjectives in responses are thought to most associate the content of an image.
+Therefore, we first perform part-of-speech (POS) tagging on the model responses and count the average number of noun/adjective tokens and other types of tokens.
+Specifically, in the ground-truth responses, 39.6% of the tokens are nouns or adjectives.
+In the responses from LLAVA-1.5-7B, the proportion of noun and adjective tokens remains nearly constant at 39.2%, both before and after TPO.
+Afterwards, we count the average score from Equation 4 of noun/adjective tokens and other types of tokens.
+The results show that noun and adjective tokens have significantly higher scores than other types, indicating higher relevance to images.
+After applying TPO, these scores of all the tokens increased notably.
+The results supports our conclusions: 1) The visual-anchored rewards reflects token-image relevance.
+2) TPO enhances the alignment of generated tokens with image content.
+Attentions
+To further validate TPO’s effectiveness in enhancing visual alignment, we measure the relevance using the sum of attention weights between responses and images.
+On the MMHal dataset, the overall image attention weights for LLaVA-1.5-7B increased from 0.14 before TPO training to 0.17 afterward.
+Additionally, Figure 3 visualizes the cases, showing a significant increase in image attention weights for response tokens, especially for visual-anchored tokens (e.g., table, cord).
+This highlights our method’s success in improving the model’s integration of visual information, thus reducing hallucinations.
+Self-Calibration
+To illustrate that our method enables the model to progressively enhance its focus on visual information through continuous self-calibration during training, we present the evolution of scores for positive and negative samples, as calculated by Equation (5), across various training steps.
 With $a=0.5$, it follows that $c_{y_{i}}\in(0.5,1.5)$.
@@
 This trend illustrates the self-calibrating effect of our method, which ultimately enhances the model’s ability to focus on visual information.
-Figure 5: The curve of changes in self-calibrated rewards for positive and negative samples over training steps, with a sample point taken every 10 steps.
 5 Conclusion
-In this study, we propose TPO, a novel token preference optimization framework, aimed at mitigating hallucinations in LVLMs.
+In this study, we propose a novel pereference alignment method, TPO, to mitigation hallucinations in LVLMs.
 TPO incorporates a self-calibrated visual-anchored reward mechanism that automatically identifies "vision-anchored tokens" and adaptively assigns appropriate rewards to them.
@@
 All the models and datasets we used are open source, so we believe that the work in this paper does not pose any potential threats.
-Appendix A Setup Details
+Appendix A Implement Details
+A.1 Setup
 In our experiments, we trained the LLaVA-v1.5 model.
 For our TPO method and the vanilla DPO method, we set the maximum learning rate to 5e-8 on the 7B version and trained for 4 epochs.
-We set the maximum learning rate to 2e-7 on the 7B version and trained for 4 epochs.
+We set the maximum learning rate to 2e-7 on the 13B version and trained for 4 epochs.
 The RLHF-V training was set according to the paper Yu et al.
 (2024b).
-All parts requiring GPT-4 evaluation use the GPT-4-0613 version, and the MM-Vet testing is conducted on the official evaluation website.
-Our experiments were all conducted on a server equipped with 8 GPUs; in specific cases (such as the 13B model), we utilized 32 GPUs.
-For the hyperparameter setting, all hyperparameters are consistent with those of our main experiment.
-Moreover, the level of diffusion noise in our model is represented by a formula $\xi=Sigmoid(l_{t})\times(0.5\times 10^{-2}-10^{-5})+10^{-5}$, where $l_{t}$ is a list of 1,000 numbers taken at equal intervals over the interval $[-6,6]$.
-The cases in Figure 1 and Figure 4 come from benchmarks Sun et al. (2023), while the cases in Figure 2 come from the RLHFV training set Yu et al. (2024b).
-Appendix B Experimental Results
-In this section, we will provide a detailed presentation of further ablation studies on the noise steps and $a$ in Eq 6, as shown in Table 4 and Table 5.
-The optimal performance for noise steps is attained at the 500-steps.
-This level of medium corruption allows the model to capture the general outline of the image while neglecting finer details, which can lead to the generation of hallucinations in the visual-anchored tokens.
-and The Optimal performance for $a$ is $0.5$, when the visual-anchored score $s=0$, we get $c=1$, not introducing additional reward signals can yield better outcomes.
+All parts requiring GPT-4 evaluation use the GPT-4-0613 8K version, and the MM-Vet testing is conducted on the official evaluation website.
+We also trained the Qwen2-vl model. For our TPO method and the vanilla DPO method, we set the maximum learning rate to 5e-9 and trained for 4 epochs.
+For a fair comparison, we set the seed to 42 during training and greedy decoding was used during inference.
+Our experiments were all conducted on a server equipped with 8 Nivdia A100 GPUs; in specific cases (such as the 13B model), we utilized 32 Nivdia A100 GPUs.
+For the hyperparameter settings, all hyperparameters are consistent with those of our main experiment.
+Moreover, the level of diffusion noise in our model is represented by a formula $\xi=Sigmoid(l_{t})\times(0.5\times 10^{-2}-10^{-5})+10^{-5}$, where $l_{t}$ is a list of 1,000 numbers taken at equal intervals over the interval $[-6,6]$, and $\epsilon\in N(0,1)$.
+The cases in Figure 1 and Figure 3 come from benchmarks Sun et al. (2023), while the cases in Figure 2 come from the RLHF-V training set Yu et al. (2024b).
+A.2 Benchmarks
+The three hallucination benchmarks: (1) AMBER : a multi-dimensional hallucination benchmark with more than 15K samples, including discriminative and description tasks.
+(2) MMHal-Bench : it measures the hallucination rate and informativeness of responses.
+(3) HallusionBench : it evaluates visual illusions and knowledge hallucinations through systematically structured discriminative tasks.
+The four general benchmarks: (1) SEED Bench : a benchmark for LVLMs on generative comprehension.
+(2) MMBench: a comprehensive benchmark designed to evaluate the capabilities across various tasks and modalities.
+(3) LLaVA Bench: a benchmark for evaluating multi-modal conversation, detailed description, and complex reasoning.
+(4) MM-Vet: a benchmark to assess integrated capabilities.
+A.3 Training Efficiency
+In TPO, generating corrupted images at each step incurs almost no time cost, as it is done during the initial data preparation.
+The main time consumption comes from calculating logits $p_{log}(y_{i}|x,v_{c},y_{<i})$ for the noisy images.
+We have also conducted a careful analysis of the time consumption for LLava-1.5-7B under the settings in Section A.1, the training durations for DPO and TPO were 1 hour 24 minutes and 1 hour 57 minutes, respectively, indicating about a 40% increase in time.
+Nevertheless, all training methods aimed at eliminating hallucinations inevitably incur additional time costs, compared to other methods requiring fine-grained annotations, our self-calibrated approach with 40% time increase proves to be sufficiently efficient.
+It has also shown superior outcomes on 5K training data training to CSR training on 13K data and POVID training on 17K data.
+This highlights the efficacy of our method in guiding the model to pay more attention to image details and in reducing hallucinations.
+We promise we will further elaborate on our efficiency in detail in the final version.
+Appendix B Ablation Analysis
+Noise Step
+We ablate on the noise steps in Figure 5 (a).
+As shown, the optimal performance is achieved at the step of 500.
+This medium corruption enables the model to grasp the general outline of the image while missing the detailed contents, which is prone to generate hallucinations of the visual-anchored tokens.
+The Figure also shows when step=0, TPO still effective and significantly better than DPO.
+This confusion is a code-implementation issue.
+In implementation as shown in Listing 1, we first convert the image into a tensor, add noise, and then convert it back into an image.
+This encode-decode process introduces some losses.
+Our method of setting the noise step to 0 serves as an ablation experiment to test the impact of this loss on our method, and it allows our experiment to more comprehensively demonstrate the advantages of TPO.
+The following portion of code may help you better understand our encode-decode process for adding noise.
+We will also open source all the code once the paper is accepted.
+Parameter $\bf a$
+We conduct experiments by varying the parameter $a$ introduced in Equation (5) with the results shown in Figure 5 (b).
+By setting $a=[0,0.5,1]$, we observed consistently good performance across all configurations.
+This suggests that effective performance is achieved as long as the reward mechanism successfully highlights token differences and identifies visually anchored tokens.
+Notably, the best overall results are obtained with $a=0.5$, validating our proposed method and hypothesis.
+This indicates that when the visual-anchored score $s=0$, setting $c=1$, not introducing additional reward signals can yield better outcomes.
+pil_to_tensor = transforms.ToTensor()
+tensor_to_pil = transforms.ToPILImage()
+image = Image.open(default_image_path).convert("RGB")
+image_tensor = pil_to_tensor(image)
+image_noisy = add_diffusion_noise(image_tensor, 500)
+image_noisy = tensor_to_pil(image_noisy)
+Listing 1: Example Python Code for Noise Addition
+Appendix C Comparison with Decoding-based Methods
+In addition to preference alignment, many researches have improved the decoding strategy of the model to eliminate hallucinations.
+In this section, we selected some typical works on decoding strategies, such as VCD Leng et al.
+(2024), OPERA Huang et al.
+(2024) and LURE Zhou et al.
+(2023), and compared the results with them.
+The results show that TPO achieves more significant hallucination reduction.
+Preference alignment and decoding strategies are two important and parallel categories of methods for hallucination mitigation.
+We believe that training with preference alignment offers several advantages: 1) Direct Optimization of Output Preferences: This approach directly optimizes the model’s output to align with desired preferences without requiring changes to the decoding strategy.
+2) Higher Inference Efficiency: Preference alignment typically results in more efficient inference, as it does not introduce additional complexity during the decoding process.
+One key advantage of decoding methods is that they do not require retraining the model, making them highly efficient for deployment.
+However, this does not preclude the benefits of preference alignment.
+In fact, we believe combining these two approaches can yield even better results.
+Appendix D Results on Object Hallucination
+In the AMBER benchmark, there is a subset for evaluating object hallucinations in image description tasks.
+Since this paper focuses on visual question answering, this part of the experiment is included in this section.
+To assess the proportion of object hallucinations in image descriptions, AMBER uses Chair as the metric.
+The results are shown in Figure 6.
+Note that ’Chair’ represents the hallucination ratio, where a smaller value indicates better model performance.
+To more clearly illustrate the comparison between methods in the figure, we use $10-chair$ as the indicator.
+The results show that TPO can not only mitigate the hallucination in visual question answering, but also eliminate the object hallucination in image descriptions to a certain extent.
+Method AMBER MMHal
+[…]
+VCD 71.8 74.9 2.12 54.20
+LURE 73.5 77.7 1.64 60.40
+OPERA 75.2 78.3 2.15 54.20
+LLaVA-1.5-7B 71.7 74.3 2.01 61.46
++DPO 77.5 82.1 2.14 58.33
++TPO 79.3 85 2.47 51.04
+Table 8: Comparison of Results
+Figure 6: Chair Performance Comparison.
+Method AMBER MMHal HallusionBench
+[…]
+LLaVA-1.5-7B 71.7 74.3 2.01 61.5 42.6 41.2 47.2
++TPO (white) 78.0 82.7 2.26 55.2 44.2 45.4 49.3
++TPO 79.3 85.0 2.5 51.0 41.8 48.4 50.2
+Table 9: Comparison of different noise adding method. “white" indicates that blank images are used in place of noisy images.
+Appendix E Comparison of Different Noise Adding Methods.
+To evaluate the impact of different methods of adding noise to images on our approach, we test a scheme where noise images were replaced with white images under the same experimental conditions.
+The results, shown in Table 9, demonstrate the superior performance of our method.
+We believe that the noise addition method used in our paper can control noise levels to create images that are more likely to induce hallucinations in the model, thereby achieving better results.

### v4

@@
 For example, by building on top of LLaVA and Qwen, our TPO boosts the performance absolute improvement for hallucination benchmarks.
+Our code is available at https://github.com/alibaba/TPO.
 Token Preference Optimization with Self-Calibrated Visual-Anchored Rewards for Hallucination Mitigation
-Jihao Guthanks: Equal contribution.thanks: Work done during an internship at Alibaba Group.1, Yingyao Wang11footnotemark: 11, Meng Cao2, Pi Bu1, Jun Songthanks: Corresponding Author.1, Yancheng He1, Shilong Li1, Bo Zheng1 1Alibaba Group 2Mohamed bin Zayed University of Artificial Intelligence {gujihao.gjh, wangyingyao.wyy, jsong.sj}@taobao.com
+Jihao Guthanks: Equal contribution.thanks: Work done during an internship at Alibaba Group.1, Yingyao Wang11footnotemark: 11, Meng Cao2, Pi Bu1, Jun Songthanks: Corresponding Author.1, Yancheng He1, Shilong Li1, Bo Zheng1 1Taobao & Tmall Group of Alibaba 2Mohamed bin Zayed University of Artificial Intelligence {gujihao.gjh, wangyingyao.wyy, jsong.sj}@taobao.com
 1 Introduction
@@
   • Extensive experiments on the popular hallucination benchmarks demonstrate the state-of-the-art performance of the proposed TPO.
-Methods Visual- Anchored Token- level Non Fine-grained Annotations
+Methods
+Visual-
+Anchored
+Token-
+level
+Non Fine-grained
+Annotations
 DPO
@@
 CSR
+MDPO
 V-DPO
@@
 (2024a), CSR Zhou et al.
-(2024b), RLHF-V Yu et al.
-(2024b), V-DPO Xie et al.
-(2024) and our proposed TPO.
+(2024b), MDPO Wang et al.
+(2024a), V-DPO Xie et al.
+(2024), RLHF-V Yu et al.
+(2024b).
 2 Related Works
@@
 (2023); Shao et al.
+(2024a); Jiang et al.
+(2024); Yue et al.
+(2024); Xiao et al.
+(2025); Sarkar et al.
+(2024); Zhao et al.
 (2023).
@@
 (2024a) method constructs negative samples for preferred data by adding noise to the image and providing hallucinated patterns to guide the model to generate hallucinated responses.
-The pre-structured data will be employed for off-policy DPO training.
+The MDPO incorporate optimization for image preference, training with the images before and after alteration as positive and negative samples.
 Apart from these works, RLAIF Yu et al.
@@
 (2024b) methods, which are built upon on-policy DPO strategy, construct preference pairs by iteratively performing self-rewarding to select preference pairs.
-A notable commonality among the aforementioned studies is that assigning response-level rewards for each generated sequence is insufficient for effectively aligning with genuinely hallucination-prone contents.
+R1-Onevision Yang et al.
+(2025) enhances the visual reasoning capabilities by employing Group Relative Policy Optimization (GRPO).
+However, assigning response-level rewards for each generated sequence is insufficient for effectively aligning with genuinely hallucination-prone contents.
 Other studies, RLHF-V Yu et al.
@@
 It is worth mentioning that CSR also considered this problem and introduced CLIP Radford et al.
-(2021) to calculate the relevance score between generated text and vision information as an additional reward.
-However, this method requires the introduction of an additional model, which reduces the training efficiency.
+(2021) to calculate the relevance score between generated text and vision information as an additional reward, and TLDR Fu et al.
+(2024) score each token by training a scoring model.
+However, these methods requires the introduction of an additional model, which reduces the training efficiency.
 In this paper, we propose a token-level preference optimization method with self-calibrated visual-anchored rewards (TPO), aimed at addressing the aforementioned challenges.
@@
 $v_{c}(k)=\sqrt{\bar{\xi}_{k}}\cdot v+\sqrt{1-\bar{\xi}_{k}}\cdot\epsilon,$ (3)
-where $\xi$ is a predefined noise parameter derived from a list with 1,000 equally spaced elements [^1] More details can be found in Appendix A, and experimental analysis can be found in Appendix E.
+where $\xi$ is a predefined noise parameter derived from a list with 1,000 equally spaced elements [^1] More details can be found in Appendix A, and experimental analysis can be found in Appendix D.
 $\bar{\xi}_{k}$ is a cumulant, i.e., $\bar{\xi}_{k}=\prod_{i=0}^{k}\xi_{i}$.
@@
 $s_{y_{i}}=p_{log}(y_{i}|x,v,y_{<i})-p_{log}(y_{i}|x,v_{c},y_{<i}),$ (4)
-where $s_{y_{i}}$ denotes the distribution difference of the token $y_{i}$ of the response $y$.
-$p_{log}$ refers to the raw logits output of the model, before applying softmax normalization.
-One example case is demonstrated in Figure 1, which demonstrates that $s$ reflects the visual relevance of each token $y_{i}$.
+where $s_{y_{i}}$ denotes the distribution difference of the token $y_{i}$ of the response $y$. $p_{log}$ refers to the raw logits output of the model, before applying softmax normalization.
+The sole variation in the input when computing $s_{y_{i}}$ of the token lies in whether noise is introduced into the image.
+A higher contrastive score, reflecting a greater divergence in logits, indicates that the current token is more susceptible to image perturbations.
+This increased sensitivity suggests a stronger dependence on and relevance to visual information, thereby marking it as a visual-anchored token.
+One example case is demonstrated in Figure 1, which demonstrates that $s_{y_{i}}$ reflects the visual relevance of each token $y_{i}$.
 Then, a self-calibration process is proposed to generate the final visual-anchored rewards $c_{y_{i}}$.
@@
 $c_{y_{i}}^{\theta}$ is continuously updated at each step during training as the model changes.
-To calculate each token in the entire reward function, we add a term $log\frac{c_{y_{i}}^{\theta}}{c_{y_{i}}^{\text{ref}}}\in(-log3,log3)$, which has a reasonable upper and lower bound.
+To calculate each token in the entire reward function, we add a term $log\frac{c_{y_{i}}^{\theta}}{c_{y_{i}}^{\text{ref}}}\in(-log3,log3)$ (as we set $a=0.5$ in Equation (5)), which has a reasonable upper and lower bound.
 For positive samples, this term is expected to increase, while for negative samples, it is expected to decrease.
@@
 Method AMBER MMHal HallusionBench General Benchmarks
-Acc F1 Score Hal $\downarrow$ Easy Hard aAcc SEED MMB LLaVA MM-Vet
+Acc F1 Score Hal $\downarrow$ Easy Hard aAcc SEED MMB LLaVA MMVet
+R1-Onevision 80.2 85.7 3.85 36.46 63.74 50.47 62.80 35.2 83.7 67.8
 LLaVA-1.5-7B 71.7 74.3 2.01 61.46 42.64 41.16 47.21 66.1 73.3 65.6 31.6
@@
 + RLHF-V 74.8 78.5 2.02 60.42 42.20 43.72 48.27 66.1 73.1 68.0 32.3
++ MDPO 2.39 54.00
 + V-DPO 81.6 2.16 56.00 51.63
@@
 + TPO (Ours) 83.9 88.0 2.72 45.83 44.40 46.05 50.93 68.7 76.8 72.8 36.2
+Qwen2-VL-7B 86.5 90.0 3.5 29.0 67.0 48.8 64.0 45.0 79.0 82.4 61.4
++ DPO 86.5 90.0 3.7 28.1 67.3 49.3 64.5 45.0 79.0 81.9 60.2
++ TPO (Ours) 86.4 89.9 4.2 18.8 67.9 50.0 65.2 45.0 79.0 82.9 61.4
 Table 2: Performence of LLaVA-1.5 on hallucination and general benchmarks.
@@
 (2024b) are based on our testing of their open-source model weights, while the results for V-DPO Xie et al.
-(2024) are taken from previous work (bold: the best score; underline: the second best)
-Method AMBER MMHal HallusionBench
-[…]
-Qwen2-vl 86.5 90.0 3.5 29.0 67.0 48.8 64.0
-+DPO 86.5 90.0 3.7 28.1 67.3 49.3 64.5
-+TPO 86.4 89.9 4.2 18.8 67.9 50.0 65.2
-Table 3: Performence of Qwen2-vl-7B on hallucination benchmarks.
+(2024), MDPO Wang et al.
+(2024a) are taken from previous work
 4 Experiment
@@
 Furthermore, to evaluate the effectiveness of TPO on more advanced and powerful model, we implement TPO training based on Qwen2-VL Wang et al.
-(2024), and compare it with the DPO method.
+(2024b), and compare it with the DPO method.
 For the dataset, we directly utilize the preference pairs provided by RLHF-V (5K) without their fine-grained human annotations.
@@
 (2024c).
-In addition, we provide the results of its Chair metric in Appendix D.
+In addition, we provide the results of its Chair metric in Appendix C.
 Moreover, we also evaluate the performance of TPO on four general benchmarks: SEED Bench Li et al.
@@
 Baselines
-We mainly compare TPO with the LLaVA-1.5-7B SFT model, as well as with the DPO and V-DPO Xie et al.
-(2024) methods trained using RLHF-V Yu et al.
+We mainly compare TPO with the R1-Onevision Yang et al.
+(2025), LLaVA-1.5-7B SFT model, as well as with the V-DPO Xie et al.
+(2024) and DPO methods trained using RLHF-V Yu et al.
 (2024b) data, along with two improved methods, CSR Zhou et al.
@@
 Additionally, to demonstrate the advantages of TPO, we reproduced the strong baseline method, RLHF-V, on LLaVA-1.5-13B and conducted a comparison.
-To further demonstrate the effectiveness of TPO on more advanced and powerful model, we additionally employ Qwen2-vl-7B as the baseline mode and compare our TPO with DPO.
+Furthermore, we additionally employ Qwen2-VL-7B Wang et al.
+(2024b) as the baseline mode and compare our TPO with DPO.
+4.2 Main Results
+[…]
+Notably, on the HallusionBench evaluation metrics, "Easy" represents the accuracy of original image-based questions, which tend to rely on prior knowledge, while "Hard" represents the accuracy of questions based on manually edited images, which tend to rely on visual information.
+Our method leads to the most significant improvement for the original model on hard questions.
+This indicates that our approach enables the model to focus more on visual information rather than textual prior knowledge to provide accurate answers.
+In general benchmarks, our approach remains stable against the backbone models and achieves improvement on most benchmarks.
+We attribute it to our method helping the model associate with more visual information when answering questions.
+This shows that our approach can improve hallucination issues while maintaining good performance in general evaluation tasks.
+4.3 Results on Qwen2-VL-7B
+As Table 2 shown, we report the results on the key metrics of three hallucination benchmarks.
+The results indicate that our TPO outperforms DPO on most benchmarks.
+On Qwen2-VL-7B, which has strong inherent capabilities, using 5K RLHF-V data for DPO alignment barely improves the performance.
+However, introducing TPO leads to a significant further enhancement.
+This demonstrates that TPO can capture and learn more subtle preferences from the data and brings higher data utilization efficiency.
+In addition, the results on the chair metric in Figure 6 further demonstrate that TPO can also significantly solve the object hallucination problem of Qwen2-VL-7B.
+Figure 3: Comparison of attention weights for LLaVA before and after TPO training.
+Each horizontal line represents the mean of that data.
+The blue section response incorrectly, with many ’visual-anchored tokens’ tokens having high attention weights but resulting in hallucinated responses (e.g.
+USB).
+The red section on the right answered correctly.
+Figure 4: The curve of changes in self-calibrated rewards for positive and negative samples over training steps, with a sample point taken every 10 steps.
+4.4 Ablation Studies
+Visual-Anchored Rewards
+Table 3 demonstrates that TPO can enhance model performance when rewards are assigned separately to positive and negative samples, achieving results comparable to those obtained by rewarding both simultaneously.
+However, by providing opposite rewards to positive and negative samples, where rewards are negatively correlated with the visual relevance of positive samples and positively correlated with that of negative samples, TPO’s performance significantly deteriorates.
+In some metrics, this approach yields even poorer results than the original LLaVA-1.5 model.
+This further underscores the validity of the designation of visual-anchored rewards.
+Noise Step
+We conduct an ablation study on the noise steps, as illustrated in Figure 5 (a) and detailed in Table 5 of Appendix.
+The results indicate that optimal performance is achieved at 500 steps.
+This intermediate level of corruption allows the model to retain the general structure of the image while obscuring fine-grained details, thereby reducing the tendency to generate hallucinations among visual-anchored tokens.
+These findings are consistent with previous work Zhou et al.
+(2024a) which noted that as image noise increases, models tend to produce responses such as “there are noise spots in the image,” while the probability of hallucination first increases and then decreases.
+Our experimental results align with this trend and provide further empirical support.
+The Figure also shows when step=0, TPO still effective and significantly better than DPO.
+This confusion is a code-implementation issue.
+In implementation as shown in Listing 1, we first convert the image into a tensor, add noise, and then convert it back into an image.
+This encode-decode process introduces some losses.
+Our method of setting the noise step to 0 serves as an ablation experiment to test the impact of this loss on our method, and it allows our experiment to more comprehensively demonstrate the advantages of TPO.
+The following portion of code may help you better understand our encode-decode process for adding noise.
+We will also open source all the code once the paper is accepted.
+Parameter $\bf a$
+We conduct experiments by varying the parameter $a$ introduced in Equation (5) with the results shown in Figure 5 (b) (detailed in Table 6 of Appendix).
+By setting $a=[0,0.5,1]$, we observed consistently good performance across all configurations.
+This suggests that effective performance is achieved as long as the reward mechanism successfully highlights token differences and identifies visually anchored tokens.
+Notably, the best overall results are obtained with $a=0.5$, validating our proposed method and hypothesis.
+This indicates that when the visual-anchored score $s=0$, setting $c=1$, not introducing additional reward signals can yield better outcomes.
+Figure 5: Performance curves with the change of the noise steps-(a) and the change of parameter aa-(b), We separately present the F1 of AMBER, the hallucination rate of MMHAL, the aACC of HallusionBench, and the average value of the general benchmarks.
 Method AMBER MMHal HallusionBench General Benchmarks
-Acc F1 Score Hal $\downarrow$ Easy Hard aAcc SEED MMB LLaVA MM-Vet
+Acc F1 Score Hal $\downarrow$ Easy Hard aAcc SEED MMB LLaVA $M$-Vet
 LLaVA-1.5-7B 71.70 74.3 2.01 61.46 42.64 41.16 47.21 66.1 73.3 65.6 31.6
@@
 Response of LLaVA and LLaVA+TPO (TPO) correspond to the outputs before and after TPO training and the scores calculated by LLaVA-1.5-7B and LLaVA-1.5-7B+TPO, corresponding to Fiure 1.
-4.2 Main Results
-[…]
-Notably, on the HallusionBench evaluation metrics, "Easy" represents the accuracy of original image-based questions, which tend to rely on prior knowledge, while "Hard" represents the accuracy of questions based on manually edited images, which tend to rely on visual information.
-Our method leads to the most significant improvement for the original model on hard questions, with a slight improvement on easy questions.
-This indicates that, compared to other methods, our approach enables the model to focus more on visual information rather than the textual prior knowledge to provide accurate answers.
-On general benchmarks, our approach remains stable against the original LLaVA model and achieves the greatest improvement on most benchmarks.
-We attribute it to that our method helps the model associate with more visual information when answering questions.
-This shows that our approach can improve hallucination issues while maintaining good performance on general evaluation tasks.
-The results also indicate that TPO introduces performance fluctuations on the LLaVABench and MMVet metrics for LLaVA-1.5-13B.
-This occurs because the training data primarily focuses on aligning preferences within the context of hallucinations, which inevitably leads to slight overfitting.
-Although augmenting the dataset with task-specific data could mitigate this issue, it is not the research focus of this paper.
-4.3 Results on Qwen2-vl
-As Table 3 shown, we report the results on the key metrics of three hallucination benchmarks.
-The results indicate that our TPO outperforms DPO on most benchmarks.
-On Qwen2-vl, which has strong inherent capabilities, using 5K RLHF-V data for DPO alignment barely improves the performance.
-However, introducing TPO leads to a significant further enhancement.
-This not only demonstrates the effectiveness of TPO on other powerful models but also shows that TPO’s token-level visual anchor rewards brings higher data utilization efficiency.
-Although Qwen2-vl shows almost no room for improvement on AMBER, TPO still achieves significant improvement on the Chair metric, as shown on Figure 5.
-TPO can capture and learn more subtle preferences from the data.
-Figure 3: Comparison of attention weights for LLaVA before and after TPO training.
-Each horizontal line represents the mean of that data.
-The blue section response incorrectly, with many ’visual-anchored tokens’ tokens having high attention weights but resulting in hallucinated responses (e.g.
-USB).
-The red section on the right answered correctly.
-Figure 4: The curve of changes in self-calibrated rewards for positive and negative samples over training steps, with a sample point taken every 10 steps.
-4.4 Ablation Studies
-Visual-Anchored Rewards
-Table 4 demonstrates that TPO can enhance model performance when rewards are assigned separately to positive and negative samples, achieving results comparable to those obtained by rewarding both simultaneously.
-However, by providing opposite rewards to positive and negative samples, where rewards are negatively correlated with the visual relevance of positive samples and positively correlated with that of negative samples, TPO’s performance significantly deteriorates.
-In some metrics, this approach yields even poorer results than the original LLaVA-1.5 model.
-This further underscores the validity of the designation of visual-anchored rewards.
-Hyperparameters
-To optimize the hyperparameters in TPO, we perform comparative experiments on the noise steps (Section 3.2) and parameter $a$ in Equation 5.
-As the Figure 5 shown, the model performs best with 500 noise steps.
-Testing $a$ across the range $a=[0,0.5,1]$, we find that $a$=0.5 produces the best outcome.
-This supports our hypothesis that setting $s=0$ and $c=1$, without adding extra reward signals, leads to superior performance [^2] The more detail results and discussions are provided in Appendix..
-Figure 5: Performance curves with the change of the noise steps-(a) and the change of parameter $a$-(b), We separately present the F1 of AMBER, the hallucination rate of MMHAL, the aACC of HallusionBench, and the average value of the general benchmarks.
-More detailed metrics can be found in the Appendix B.
 4.5 Analysis
@@
 By adding noise to the visual input and capturing changes in the generation probability of each token, TPO computes a score indicating each token’s relevance to visual information.
-Subsequently, a self-calibration process adjusts these scores to ensure that: 1) Rewards for positive examples exceed those for negative examples.
-2) The final rewards enhance the focus on visual information for all of "vision-anchored tokens" in preference data.
 Based on the self-calibrated visual-anchored reward, TPO can perform more efficient token-level preference alignment optimization for LVLMs.
@@
 It can enable the model to specifically improve its focus on image information in certain domains, thus having extensive industrial applications.
+Besides, we believe that the core part of the TPO method, the visual-anchored reward scoring method, possesses strong extensibility.
+For example, these token-level rewards can also be used to weight the probability distributions in the calculations for other RLHF methods, enhancing the visual attention of multimodal models.
 We will continue to expand in this direction, and we believe that the technology we have proposed in this paper has a vast space for further development and application.
@@
 All parts requiring GPT-4 evaluation use the GPT-4-0613 8K version, and the MM-Vet testing is conducted on the official evaluation website.
-We also trained the Qwen2-vl model. For our TPO method and the vanilla DPO method, we set the maximum learning rate to 5e-9 and trained for 4 epochs.
+During the training process, we froze the vision encoder and only trained the LLM.
+We also trained the Qwen2-VL model. For our TPO method and the vanilla DPO method, we set the maximum learning rate to 5e-9 for 7B model, 1e-9 for 2B model and trained for 4 epochs.
 For a fair comparison, we set the seed to 42 during training and greedy decoding was used during inference.
@@
 For the hyperparameter settings, all hyperparameters are consistent with those of our main experiment.
-Moreover, the level of diffusion noise in our model is represented by a formula $\xi=Sigmoid(l_{t})\times(0.5\times 10^{-2}-10^{-5})+10^{-5}$, where $l_{t}$ is a list of 1,000 numbers taken at equal intervals over the interval $[-6,6]$, and $\epsilon\in N(0,1)$.
+We used ‘Pytorch’ in our code.
+Moreover, the level of diffusion noise in our model is represented by a formula $\xi=Sigmoid(l_{t})\times(0.5\times 10^{-2}-10^{-5})+10^{-5}$, where $l_{t}$ is a list of 1,000 numbers taken at equal intervals over the interval [6,6][-6,6], and $\epsilon\in N(0,1)$.
 The cases in Figure 1 and Figure 3 come from benchmarks Sun et al. (2023), while the cases in Figure 2 come from the RLHF-V training set Yu et al. (2024b).
@@
 This highlights the efficacy of our method in guiding the model to pay more attention to image details and in reducing hallucinations.
-We promise we will further elaborate on our efficiency in detail in the final version.
-Appendix B Ablation Analysis
-Noise Step
-We ablate on the noise steps in Figure 5 (a).
-As shown, the optimal performance is achieved at the step of 500.
-This medium corruption enables the model to grasp the general outline of the image while missing the detailed contents, which is prone to generate hallucinations of the visual-anchored tokens.
-The Figure also shows when step=0, TPO still effective and significantly better than DPO.
-This confusion is a code-implementation issue.
-In implementation as shown in Listing 1, we first convert the image into a tensor, add noise, and then convert it back into an image.
-This encode-decode process introduces some losses.
-Our method of setting the noise step to 0 serves as an ablation experiment to test the impact of this loss on our method, and it allows our experiment to more comprehensively demonstrate the advantages of TPO.
-The following portion of code may help you better understand our encode-decode process for adding noise.
-We will also open source all the code once the paper is accepted.
-Parameter $\bf a$
-We conduct experiments by varying the parameter $a$ introduced in Equation (5) with the results shown in Figure 5 (b).
-By setting $a=[0,0.5,1]$, we observed consistently good performance across all configurations.
-This suggests that effective performance is achieved as long as the reward mechanism successfully highlights token differences and identifies visually anchored tokens.
-Notably, the best overall results are obtained with $a=0.5$, validating our proposed method and hypothesis.
-This indicates that when the visual-anchored score $s=0$, setting $c=1$, not introducing additional reward signals can yield better outcomes.
 pil_to_tensor = transforms.ToTensor()
@@
 Listing 1: Example Python Code for Noise Addition
-Appendix C Comparison with Decoding-based Methods
-In addition to preference alignment, many researches have improved the decoding strategy of the model to eliminate hallucinations.
-In this section, we selected some typical works on decoding strategies, such as VCD Leng et al.
-(2024), OPERA Huang et al.
-(2024) and LURE Zhou et al.
-(2023), and compared the results with them.
+Appendix B Comparison with Related Methods
+To more comprehensively highlight the advantages of the TPO method, we conducted comparisons with other related works Jiang et al.
+(2024); Yue et al.
+(2024); Xiao et al.
+(2025); Sarkar et al.
+(2024); Zhao et al.
+(2023); Leng et al.
+(2024); Huang et al.
+(2024); Zhou et al.
+(2023) aimed at addressing the hallucination problem.
 The results show that TPO achieves more significant hallucination reduction.
@@
 In fact, we believe combining these two approaches can yield even better results.
-Appendix D Results on Object Hallucination
+To further elucidate the innovations of our work, we provide a more in-depth comparison with similar methods Cui et al.
+(2024); Zhou et al.
+(2024a); Leng et al.
+(2024).
+Our approach exhibits both conceptual distinctions and methodological advancements.
+Unlike FiSAO, which relies on a vision encoder to compute token-level scores, TPO derives them directly from the LVLM’s logits difference between clean and noisy images.
+Moreover, while FiSAO builds upon PPO, TPO enhances the DPO framework.
+Our method shares with VCD the idea of using noise-induced logits change to identify visual-sensitive tokens; however, TPO goes beyond identification by introducing a self-calibrating reward Eq.
+5 that jointly reinforces positive outputs, suppresses hallucinations, and strengthens image–token alignment—aiming fundamentally to improve perceptual faithfulness.
+In contrast, VCD focuses on decoding less image-affected tokens without enhancing underlying visual grounding.
+Similarly, although POVID also uses Gaussian noise for perturbation, it primarily constructs negative samples for DPO training, whereas TPO leverages noise to quantify token–image relevance and enables a more efficient token-level reinforcement learning strategy, leading to superior performance.
+Appendix C Results on Object Hallucination
 In the AMBER benchmark, there is a subset for evaluating object hallucinations in image description tasks.
@@
 Note that ’Chair’ represents the hallucination ratio, where a smaller value indicates better model performance.
-To more clearly illustrate the comparison between methods in the figure, we use $10-chair$ as the indicator.
+To more clearly illustrate the comparison between methods in the figure, we use 10chair10-chair as the indicator.
 The results show that TPO can not only mitigate the hallucination in visual question answering, but also eliminate the object hallucination in image descriptions to a certain extent.
@@
 […]
+LLaVA-1.5-7B 71.7 74.3 2.01 61.46
 VCD 71.8 74.9 2.12 54.20
@@
 OPERA 75.2 78.3 2.15 54.20
-LLaVA-1.5-7B 71.7 74.3 2.01 61.46
-+DPO 77.5 82.1 2.14 58.33
-+TPO 79.3 85 2.47 51.04
+HACL 2.13 50 - -
+EOS 2.03 59 - -
+HA-DPO 1.97 60 75.2 79.9
+HALVA 2.25 54 - 83.4
+DPO 77.5 82.1 2.14 58.33
+TPO 79.3 85 2.47 51.04
+LLAVA-1.5-13B 2.38 53 71.3 73.1
+HSA-DPO 2.61 48 - -
+HALVA 2.58 45 - 86.5
+DPO 2.47 51 83.2 86.9
+TPO 2.72 46 83.9 88
 Table 7: Comparison of Results
@@
 Table 8: Comparison of different noise adding method. “white" indicates that blank images are used in place of noisy images.
-Appendix E Comparison of Different Noise Adding Methods.
+Appendix D Comparison of Different Noise Adding Methods.
 To evaluate the impact of different methods of adding noise to images on our approach, we test a scheme where noise images were replaced with white images under the same experimental conditions.
@@
 We believe that the noise addition method used in our paper can control noise levels to create images that are more likely to induce hallucinations in the model, thereby achieving better results.
+Appendix E Evaluation of TPO Training on Real-World Performance
+To validate the effect of TPO training, we compared the performance of the LLaVA-1.5-7B models before and after training on a real-world image set of 10 images, as suggested during the review process.
+A manual analysis of the outputs revealed that the TPO-trained model provided superior results in 40% of the cases.
+In contrast, the original model was better in only 20% of cases, with both models performing equally well in the remaining 40%.
+This represents a clear net performance gain attributable to the TPO training procedure.
+Method AMBER
+Acc F1
+LLaVA-1.5-7B 71.7 74.3
++ DPO 77.5 82.1
++ CSR 73.2 76.1
++ POVID 71.9 74.7
++ RLHF-V 74.8 78.5
++ TPO (1K) 72.5 75.3
++ TPO (3K) 78.9 83.4
++ TPO (5K) 79.3 85.0
+Table 9: Evaluation of TPO Training with varying amounts of training data.
+Appendix F Analysis of TPO Training on less training data
+In order to further investigate the impact of training data volume, we constructed subsets of 1K and 3K samples from the original 5K training set.
+Using these subsets, we fine-tuned the LLaVA-1.5-7B model with a consistent learning rate of 5e-8 for 4 epochs, maintaining identical experimental settings as in the full-data scenario.
+As anticipated in table 9, the performance of the TPO method exhibits a gradual decline with reduced training data.
+Nevertheless, even when trained on only 1K samples, TPO remains competitive and surpasses most baseline methods.
+With 3K training samples, it achieves the second-highest performance, further underscoring the effectiveness and data efficiency of our approach.